### v1

# Understanding Knowledge Hijack Mechanism in In-context Learning through Associative Memory

Abstract
In-context learning (ICL) enables large language models (LLMs) to adapt to new tasks without fine-tuning by leveraging contextual information provided within a prompt. However, ICL relies not only on contextual clues but also on the global knowledge acquired during pretraining for the next token prediction. Analyzing this process has been challenging due to the complex computational circuitry of LLMs. This paper investigates the balance between in-context information and pretrained bigram knowledge in token prediction, focusing on the induction head mechanism, a key component in ICL. Leveraging the fact that a two-layer transformer can implement the induction head mechanism with associative memories, we theoretically analyze the logits when a two-layer transformer is given prompts generated by a bigram model. In the experiments, we design specific prompts to evaluate whether the outputs of a two-layer transformer align with the theoretical results.
1 Introduction
In recent years, transformer-based models, such as BERT [13] and GPT [42, 43, 5], have achieved remarkable success in natural language processing. Especially, in-context learning (ICL) [14] has emerged as a groundbreaking capability within large language models (LLMs), enabling them to adapt to new tasks without traditional fine-tuning. Instead, these models leverage patterns from a prompt or input sequence, effectively learning "in context" by interpreting examples or instructions provided in real-time [33, 56]. The capability for ICL [5] in language models develops gradually during a "phase change" observed in the early stages of training. At the same time, models acquire the ability to complete complex patterns through induction heads, suggesting that induction heads are a key mechanistic basis of ICL [40].
In the context of ICL, Jiang et al. [25] discovered a phenomenon called context hijacking. This phenomenon occurs when altering the context disrupts fact recall, causing the model to generate incorrect outputs influenced by the in-context knowledge provided in the prompt. Similarly, ICL can exhibit oversight of in-context knowledge provided in the prompt, where the model relies solely on the knowledge acquired during pretraining (global knowledge) while ignoring the in-context knowledge. These types of knowledge hijacking are thought to arise from either discarding part of the prompt’s information or overly trusting the information in the prompt. Conversely, if the model can (i) comprehensively use the information in the prompt and (ii) balance in-context knowledge and global knowledge, hijacking can be prevented. In relation to the induction head, we demonstrate that employing relative positional encoding (RPE) allows the transformer to comprehensively use the information in the prompt. Furthermore, we show that even a three-layer transformer without positional encoding can achieve comprehensive use of information. Understanding how to avoid knowledge hijacking is essential for the safe and reliable use of ICL capabilities in LLMs and is a matter of significant societal importance.
In this paper, we theoretically and experimentally carry out evaluation of oversight of in-context information in a transformer, and the model’s prioritization between in-context knowledge, derived from information within the context or the prompt, and global knowledge from the training data generated from a bigram model with triggered transitions. To summarize, we make the following contributions.
  • We theoretically show that the induction head learned by a transformer using RPE can avoid oversight of in-context knowledge within the framework of associative memory.
  • We experimentally confirmed that the model with RPE can comprehensively use in-context knowledge for the bigram model. We also demonstrated its ability to effectively infer on sequences where the transformer using absolute positional encoding (APE) fails in next-token prediction.
  • We investigated how the model prioritizes in-context knowledge versus global knowledge when generating outputs, using both theoretical analysis and experimental evaluation.
2 Related work
Paper Model Objective Theoretical analysis
Cabannes et al. [6] linear model scaling laws yes
Cabannes et al. [7] linear model learning dynamics yes
Jiang et al. [25] one-layer transformer concept association yes
Chen et al. [8] linear model noisy classification yes
Bietti et al. [4] two-layer transformer learning dynamics yes
Meng et al. [35] GPT knowledge editing no
OURS two-layer transformer
length generalization
global v.s. in-context
yes
Table 1: Comparisons among related work and ours in terms of associative memory. The concept of associative memory is widely adopted for theoretical analyses of various phenomena.
Associative memory
Associative memory refers to the storage and retrieval of patterns, and its computational model was designed to retrieve complete memories from sufficient fragments of information in neuroscience [20, 21]. After the emergence of modern Hopfield networks [30], which store patterns in an energy function and retrieve them using an updating rule. Moreover, many studies employ the concept of associative memory in various ways. For instance, Zhang et al. [58] propose a memory unit that estimates a conditional probability distribution with Gaussian kernel smoothing. Some work considers attention blocks as memory because they read and write information flowing through the residual stream [39, 17]. Meng et al. [35] locate factual association patterns in the weight matrix of an MLP. Recently, it has become increasingly popular to regard a matrix as associative memory from a theoretical perspective, as shown in Tab. 1. Previous studies examine scaling laws in relation to model capacity [6], learning dynamics of linear model in terms of particle system [7], and the effect of low-rank approximation of weight matrix to classification involving a noisy token [8] for a linear model. Additionally, Jiang et al. [25] demonstrate that a one-layer transformer can recover latent concept from long enough context. Most relevant to our research is the work by Bietti et al. [4], which mainly discusses the learning dynamics of induction heads for a two-layer transformer, but our primary focus is on the reasoning process.
In-context learning and induction head
Since Brown [5] introduced ICL, theoretical understanding of this phenomenon has been a major interest within the community. For example, Akyürek et al. [2] showed that transformer can implement a learning algorithm that solves linear regression. Von Oswald et al. [51] demonstrated that linear self-attention layer can emulate gradient descent. Similarly, many studies [1, 34, 59, 11] examined the theoretical connections between ICL and gradient descent. Mechanistic interpretability, which aims to understand the model behavior by reverse-engineering its internal components [37, 53, 10], is another perspective for analyzing ICL [40, 16]. In particular, Elhage et al. [16] discovered that two-layer attention-only transformers understand the previous use of the token and look for similar situations in the context. In addition, Olsson et al. [40] provided empirical evidence that induction heads are the mechanism behind ICL. Moreover, Bansal et al. [3] identified that many induction heads are also among the attention heads deemed critical for ICL. The most relevant study to ours is by Bietti et al. [4], who investigated the learning dynamics of the induction head mechanism in two-layer transformers from the perspective of associative memory. In contrast, our study focuses on the length generalization of the induction head mechanism and the influence of knowledge acquired during training and that obtained in context on the final output.
3 Preliminaries
3.1 Transformer architecture
A decoder-only transformer consists of a stack of identical layers, each composed of several sub-components that enable it to generate sequential outputs in an autoregressive manner. The decoder generates sequences by predicting one token at a time, conditioned on the previously generated tokens. This process continues until an end-of-sequence token is generated or the maximum sequence length is reached. Large language models commonly adopt this kind of architecture and are trained with next token prediction. In this work, we focus on a two-layer transformer architecture analyzed by Bietti et al. [4], while we use simple relative positional encoding instead of absolute positional encoding.
Calculation
Given a sequence of tokens $z_{1:T}$ of length $T\in\mathbb{N}$ from the vocabulary set $\mathcal{V}$, each token $z_{t}$ is embedded into a $d$-dimensional vector and the calculation proceeds as follows:
$x^{(0)}_{t}$ $:=w_{E}(z_{t})\in\mathbb{R}^{d},$
$x^{(1)}_{t}$ $:=\Phi_{1}x^{(0,v)}_{1:t}\sigma((W_{K}^{1}x^{(0,k)}_{1:t})^{\top}W_{Q}^{1}x^{(0,q)}_{t})$
$\qquad+x^{(0,v)}_{t}\in\mathbb{R}^{d},$ (1)
$x^{(2)}_{t}$ $:=W_{O}^{2}W_{V}^{2}x^{(1)}_{1:t}\sigma((W_{K}^{2}x_{1:t}^{(1)})^{\top}W_{Q}^{2}x_{t}^{(1)})$
$\qquad+x^{(1)}_{t}\in\mathbb{R}^{d},$ (2)
$x_{t}$ $:=W_{F}x^{(2)}_{t}+x^{(2)}_{t},$ (3)
$\hat{z}_{t+1}$ $=\operatorname*{arg\,max}_{v\in\mathcal{V}}\sigma(W_{U}x_{t})_{v},$ (4)
where $W_{E}\in\mathbb{R}^{d\times|\mathcal{V}|}$ is an embedding matrix, $\sigma:\mathbb{R}^{t}\to\mathbb{R}^{t}$ is the softmax function, $x^{(0,k)},x^{(0,q)},\text{ and }x^{(0,v)}$ are the token embeddings, which may or may not include positional information, $\Phi_{1}\in\mathbb{R}^{d\times d}$ represents $W_{O}^{1}W_{V}^{1}$, $W_{F}\in\mathbb{R}^{d\times d}$ is a linear projection, $W_{U}=(w_{U}(v_{1}),\dots,w_{U}(v_{|\mathcal{V}|}))^{\top}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is an unembedding matrix, and $v_{i}$ represents the $i$-th vocabulary in $\mathcal{V}$.
Note that Bietti et al. [4] incorporate absolute positional encoding:
$x^{(0,k)}_{t}=x^{(0,q)}_{t}=x^{(0,v)}_{t}=w_{E}(z_{t})+p_{t},$
where $p_{t}$ is the $t$-th absolute positional encoding. In contrast, we use a simplified relative positional encoding:
$x^{(0,k)}_{1:t}$ $=w_{E}(z_{1:t})+R_{1-t:0},$
$x^{(0,q)}_{t}$ $=x^{(0,v)}_{t}=w_{E}(z_{t}),$
where positional information $R_{1-t:0}=(r_{1-t},\dots,r_{-1},r_{0})\in\mathbb{R}^{d\times t}$ and $r_{-i}$ represents the positional information of the $i$-th previous token $z_{t-i}$ from the current one $z_{t}$.
3.2 Bigram model with triggered transitions
This section describes the bigram model with triggered tokens used in Bietti et al. [4]. In this bigram language model, transitions between certain "trigger" tokens (denoted as $q_{k}$) are modified in each sequence. Specifically, these trigger tokens are always followed by specific output tokens $o_{k}$. The model operates over a vocabulary of size $V$, with the following distributions: $\pi_{b}(\cdot\mid i)$ for bigram conditionals, $\pi_{u}$ for the unigram distribution, and $\pi_{o}(\cdot\mid i)$ to generate output tokens, along with a distribution $\pi_{q}$ for trigger tokens.
The process involves randomly sampling trigger tokens from $\pi_{q}$. Then, output tokens $o_{k}$ are sampled from $\pi_{o}$, and the sequence $z_{1:T}$ is generated as follows:
  • Sample $q_{1},\dots,q_{K}\sim\pi_{q}$, without replacement (random triggers).
  • Sample $o_{k}\sim\pi_{o}(\cdot\mid q_{k})$, with replacement.
  • Generate $z_{1}\sim\pi_{u}$ and for each subsequent token $z_{t}$ (for $t=2,\dots,T$), sample $z_{t}\mid z_{t-1}$ as follows:
    $p_{n}(j\mid i)=\begin{cases}\pi_{b}(j\mid i),&\text{if }i\notin\{q_{k}\},\\ 1\{j=o_{k}\},&\text{if }i=q_{k}.\end{cases}$
Motivation for the model
The induction head is a computational mechanism capable of detecting repeated patterns in context and facilitating the output of the tokens that complete these patterns when a trigger token appears. In written texts centered on a particular theme, it is common for identical patterns of tokens to recur multiple times. For instance, in texts about Harry Potter or machine learning, the token "Avada" is often followed by "Kedavra," and "neural" is typically followed by "network."
In a pure bigram model, the token following a given token $z$ is determined by conditional probabilities. As a result, consistent repeated patterns may not emerge, or multiple distinct patterns beginning with $z$ may appear, making it difficult to thoroughly evaluate the function of the induction head. However, by associating specific output tokens with designated trigger tokens, it becomes possible to generate multiple instances of the same pattern within the input sequence. Thus, this approach enables a more effective analysis of the induction head mechanism in scenarios that closely resemble real-world text.
3.3 Associative memory
Associative memory plays an important role in analyzing the behavior of memory recall at induction head in the next section. Before defining associative memory, we first adopt a technical assumption commonly used in theoretical studies on transformers, that is, that embeddings are high-dimensional random vectors, allowing them to be nearly orthogonal. Mathematically, we impose the following assumption.
Assumption 1 (Near-orthogonality [22, 31, 4]).
The embeddings $(u_{i})_{i}$ are $d$-dimensional vectors with Gaussian entries, each having a unit norm $\|u_{i}\|=1$ and $u_{i}^{\top}u_{j}=0$ for $i\neq j$. Also, $W_{0}u_{i}$ forms a new vector that has a near-unit norm and near-orthogonal to $(u_{i})_{i}$ if $W_{0}\in\mathbb{R}^{d\times d}$ is a Gaussian random matrix with appropriate entries.
With the assumption, we define associative memory.
Definition 1 (Associative Memory).
Suppose $(u_{i})_{i}$ and $(v_{j})_{j}$ are the embeddings satisfying Assumption 1. Then, A matrix is called associative memory when $W$ is expressed as the sum of the products of $u_{i}$ and $v_{j}$:
$W=\sum_{i,j}\alpha_{i,j}u_{i}v_{j}^{\top},$ (5)
where $\alpha_{i,j}\in\mathbb{R}$ is the score for the pair $(u_{i},v_{j})$.
Thanks to orthogonality, we can derive the score $\alpha_{i,j}$ for the pair $(u_{i},v_{j})$ using the operation $u_{i}^{\top}Wv_{j}$. Henceforth, we consider the weight matrices in the attention layer and the feed-forward network of the transformer as matrices that implement associative memory.
3.4 Induction head
Figure 1: The visualization of induction head mechanism.
Induction heads are a specific type of attention mechanism observed in transformer models, responsible for detecting patterns in token sequences and utilizing previous occurrences of the same token to improve predictions. The behavior of induction heads is particularly relevant for tasks that require understanding and leveraging repeated patterns, such as in language modeling.
In Fig. 1, we summarize the process by which the induction head identifies and outputs repetitive patterns in the input. When the prompt $ABC\dots A$ is given, the first-layer attention uses $W_{K}^{1}$ to copy information from the previous token to the current token. Consequently, the token $B$ retains the information of $A$ in the form of $\Phi_{1}A$. Then, in the second-layer attention, the associative memory $W_{K}^{2}$ matches the pair $\Phi_{1}A$ with $A$, attending to the position of token $B$. As a result, the information of $B$ is stored in the last token $A$ as $W_{V}^{2}B$. Finally, the output matrix $W_{O}^{2}$ transforms $W_{V}^{2}B$ into $W_{U}(B)$, producing an output that follows the repetitive pattern.
4 Theoretical result
We analyze two key aspects of induction heads: the impact of positional encoding on the oversight, and the contributions of in-context information and pretrained knowledge during inference.
4.1 Oversight of in-context knowledge
In this section, we address the theoretical aspects of oversight mechanism in the induction head mechanism. In short, for transformers with APE, it can be noted that learning of the previous token head fails, resulting in prediction errors. This issue arises not only with input sequences of lengths that were encountered during training, but also with longer sequences. Formally, we observe the representation of the weight matrix $\tilde{W}_{K}^{1}$ when weight matrices are sequentially trained, in the order of $W_{O}^{2}$, $W_{K}^{2}$, and $W_{K}^{1}$, through one step of gradient descent. We refer to Theorem 3 from Bietti et al. [4]:
Theorem 1 (Theorem 3, Bietti et al. [4]).
For any $t>1$, the learned matrix achieves the following associative memory:
$p_{t-1}^{\top}\tilde{W}_{K}^{1}p_{t}$
$\approx\frac{\eta\alpha\alpha^{\prime}(T-1)}{T^{2}t}\left\{P(t_{q}=t-1)\left(1-\frac{1}{t}\right)\right.$
$\hskip 90.00014pt\left.+O\left(\frac{1}{V}\right)\right\},$ (6)
where $p_{t}$ is the absolute positional encoding for token position $t$.
From Eq. 1, we can see that the coefficient of $p_{t-1}p_{t}^{\top}$ which constitutes the associative memory $\tilde{W}_{K}^{1}$ is inversely proportional to $t$. Therefore, as $t$ increases – meaning as we move toward the latter part of the sequence– the function of the associative memory diminishes.
In contrast, we show that transformers employing RPE consistently direct the previous-token head to attend to the preceding token, regardless of the input sequence length. We first calculate how the transformer’s weight matrices are represented by following the same learning procedure as when using a transformer with APE. Here, we present the training outcome of the matrix $W_{K}^{1}$ as associative memory, which plays the most significant role in length generalization by constructing the previous token head. We provide the complete statement and proof in Appendix B.
Theorem 2 (informal).
Under the setup described in Sec. B.2, a two-layer attention-only transformer with relative positional encoding learns the associations by one step of gradient descent for any vocabulary $v\in\mathcal{V}$:
$r_{-1}^{\top}W_{K}^{1}w_{E}(v)$
$=\frac{\eta\alpha\alpha^{\prime}}{T}\sum_{t=1}^{T}\frac{P(t_{q}=t-1)}{t}\cdot O\left(1\right)$
$\quad+\frac{\eta\alpha\alpha^{\prime}}{T}\sum_{t=1}^{T}\frac{P(t_{q}=t)}{t^{2}}\left\{O\left(\frac{1}{V}\right)+O\left(\frac{1}{T}\right)\right\},$
where $\eta\in\mathbb{R}$ is the learning rate, $\alpha,\alpha^{\prime}\in\mathbb{R}$ are constants, $t_{q}$ and $T$ are the first and second occurrence positions of trigger token, and $V$ is the size of the vocabulary set $\mathcal{V}$.
Theorem 2 suggests that the association between $r_{-1}$ and $w_{E}(v)$ is independent of both the vocabulary $v$ and the token position $t$. Therefore, regardless of the input sequence length, the previous token head will always attend to previous tokens with consistent strength.
Remark 1.
A transformer using RPE learns to attend to the previous token regardless of its position within the sequence. However, if the input sequence contains out-of-distribution tokens, the induction head mechanism does not activate. In such cases, only a transformer with APE can effectively attend to the previous token.
We also present the following Proposition that shows a three-layer transformer without positional encoding can implement the induction head mechanism.
Proposition 1.
A construction exists for a three-layer transformer without positional encoding that successfully achieves the induction head mechanism.
4.2 Global knowledge v.s. in-context knowledge
Here, we discuss whether to prioritize statistical information obtained from the training data or the information provided in-context. By calculating the logits that determine the output probabilities of the vocabulary in an associative memory transformer, we clarify the mechanisms through which outputs are generated.
Associative memory construction achieving induction head
We first introduce the following lemma, which states that the induction head mechanism is achievable by setting the appropriate weights for the matrices in a two-layer transformer with APE.
Lemma 1 (Bietti et al. [4]).
Define $U_{v}=\{u\in\mathcal{V}\mid\pi_{b}(u\mid v)>0\}$ as the support of $\pi_{b}(u\mid v)$ and $Q=\operatorname{supp}(\pi_{q})$. The induction head can be achieved by constructing matrices in the following manner.
$W_{Q}^{1}$ $=W_{Q}^{2}=I,$
$W_{K}^{1}$ $=\sum_{t=2}^{T}p_{t}p_{t-1}^{\top},$
$W_{K}^{2}$ $=\sum_{k\in Q}w_{E}(k)(W_{O}^{1}W_{V}^{1}w_{E}(k))^{\top},$
$W_{O}^{2}$ $=\sum_{v=1}^{V}w_{U}(v)(W_{V}^{2}w_{E}(v))^{\top},$
where $I$ is the identity matrix. The other matrices $W_{O}^{1}$, $W_{V}^{1}$, and $W_{O}^{2}$ are a Gaussian random matrix.
It is straightforward to prove that a two-layer transformer with RPE also has associative memory construction that achieves induction head by modifying $W_{K}^{1}=\sum_{k\in Q}w_{E}(k)r_{-1}^{\top}$. Now, using the transformer with RPE architecture, we define an associative memory transformer : a model that incorporates an induction head and a feed-forward network storing knowledge learned through pretraining.
Definition 2 (associative memory transformer).
A two-layer transformer with RPE is called associative memory transformer if its weight matrices are set as in Lemma 1 except for
$W_{K}^{1}$ $=\sum_{k\in Q}w_{E}(k)r_{-1}^{\top},$
$W_{F}$ $=\sum_{v=1}^{V}\sum_{u=1}^{V}\log{\pi_{b}(u\mid v)}w_{U}(u)w_{E}(v)^{\top},$
Note that when the probability $\pi_{b}(u\mid v)$ approaches $+0$, its logarithm $\log{\pi_{b}(u\mid v)}$ diverges to $\infty$- ∞. If we admit $\log{\pi_{b}(u\mid v)}=-\infty$ for $u\notin U_{v}$ in the construction of $W_{F}$, the transformer does not output $u$ after the token $v$ regardless of the context. This does not align with the actual behavior of the induction head. Also, due to the limitations of floating-point representation in computers, it is not possible to represent $\infty$ directly. Thus, we consider $\epsilon$ as threshold and set $\pi_{b}(u\mid v)=\epsilon$.
Associative memory transformer on limited input sequences.
First, we present the following proposition, which focuses on two token patterns in the input sequence.
Proposition 2.
Suppose a two-layer transformer is an associative memory transformer. Given a length-$T$ sequence $z_{1:T}$ where the last token is $z_{T}=q$, let the $t_{1}$-th token be $z_{t_{1}}=v_{1}$ following $z_{t_{1}-1}=q$, and let the $t_{2}$-th token be $z_{t_{2}}=v_{2}$ following $z_{t_{2}-1}=q$, where $v_{1}\neq v_{2}$ and $(3\leq)t_{1}<t_{2}$ without loss of generality. Also assume that there exists only one occurrence for each of $v_{1}$ and $v_{2}$ in the context $z_{1:T}$. Then, the logits $\xi_{v_{1}}$ and $\xi_{v_{2}}$ can be expressed as follows:
$\xi_{v_{1}}$ $=\exp{\left(\frac{e}{t_{1}+e-1}\right)}/Z+\log{\pi_{b}(v_{1}\mid q)},$ (7)
$\xi_{v_{2}}$ $=\exp{\left(\frac{1+e}{t_{2}+e-1}\right)}/Z+\log{\pi_{b}(v_{2}\mid q)},$ (8)
where $Z$ is the normalization factor of the softmax function.
We can observe a general trend from Eqs. 7 and 8. First, if the model contains global knowledge of both tokens $v_{1}$ and $v_{2}$, the output token is influenced by the knowledge, in addition to the in-context knowledge. This implies that the occurrence of tokens that rarely appear in the training dataset is suppressed. In our associative memory transformer model, global knowledge never promotes the occurrence of tokens but only suppresses inappropriate occurrence, because we always have $\log{\pi_{b}(\cdot\mid q)}\leq 0$. Second, if token patterns, such as $q\,v_{1}$ and $q\,v_{2}$, do not appear in the training data, global knowledge affects the token logits to the same extent, due to the way $W_{F}$ is defined using the threshold $\epsilon$ in Def. 2. This indicates that the model intensively suppresses token patterns that were not seen during training.
It is worth noting that an associative memory transformer exhibits a significant difference in scale when handling global knowledge and in-context knowledge. For instance, if $\pi_{b}(v\mid q)=0.1$, then $\log{\pi_{b}(v\mid q)}\approx-2.3$. Even with the functionality of an induction head, it is unlikely to predict the token patterns from the context. This is avoidable by reformulating $W_{O}^{2}$ as in Eq. 12 and adjusting the strength of induction head. We consider this setting in the next paragraph, along with more general input sequences.
If neither token appears during training, the difference between the two logits before applying the softmax function $\xi^{\prime}_{v_{2}}-\xi^{\prime}_{v_{1}}$ is calculated as follows:
$\xi^{\prime}_{v_{2}}-\xi^{\prime}_{v_{1}}=$ $\frac{e(t_{1}-t_{2})+t_{1}+e-1}{(t_{1}+e-1)(t_{2}+e-1)}$ (9)
The result implies that the token $v_{2}$ is more likely to be generated if $t_{1}$ is large and the distance between the tokens $v_{1}$ and $v_{2}$, i.e., $t_{2}-t_{1}$ is small. This trend arises from the first layer of the attention block, where the attention is not fully concentrated on the previous token, resulting in a diffusion of attention across other tokens as well. In other words, the more tokens that precede the current token, the lower the attention score assigned to the current token.
Stronger associative memory on general input sequences.
Next, we introduce a more sophisticated analysis of in-context and global knowledge by relaxing assumptions in the sequence $z_{1:T}$. The transformer considered here has associative memory with larger scores, characterized by a weighted sum of the individual terms of the weight matrices in Sec. 3.4, where each term is scaled by an appropriate coefficient. We consider a special case in which each term has the same coefficient $\tau_{1},\tau_{2},\tau_{3}\in\mathbb{R}$.
Definition 3 (stronger associative memory transformer).
A stronger associative memory transformer is equipped with the following weight matrices:
$W_{K}^{1}$ $=\tau_{1}\sum_{k\in Q}w_{E}(k)r_{-1}^{\top},$ (10)
$W_{K}^{2}$ $=\tau_{2}\sum_{k\in Q}w_{E}(k)(W_{O}^{1}W_{V}^{1}w_{E}(k))^{\top},$ (11)
$W_{O}^{2}$ $=\tau_{3}\sum_{v=1}^{V}w_{U}(v)(W_{V}^{2}w_{E}(v))^{\top},$ (12)
with the parameters $\tau_{1},\tau_{2},\tau_{3}\in\mathbb{R}_{+}$. The other weight matrices remain the same.
Proposition 3.
Suppose a two-layer transformer is a stronger associative memory transformer as defined in Eqs. 10, 11, and 12. Given a length-$T$ sequence $z_{1:T}$ where the last token is $z_{T}=q$, let $f(v)$ be the number of token pattern "$q\,v$" appearing in the sequence for vocabulary $v\in\mathcal{V}$. For sufficiently large $\tau_{1}$ and $\tau_{2}$, the logits $\xi_{v}$ can be expressed as follows:
$\xi_{v}$ $\underset{\begin{subarray}{c}\tau_{1}\\ \tau_{2}\end{subarray}}{\approx}\log{\pi_{b}(v\mid q)}$
$\quad+\tau_{3}\cdot\frac{f(v)+\mathbf{1}\{v=q\}\mathbf{1}\{z_{1}=q\}}{\left(\sum_{v^{\prime}=1}^{V}f(v^{\prime})\right)+\mathbf{1}\{z_{1}=q\}}.$ (13)
It can be observed that the transformer equipped with a general associative memory is affected by the proportions of each token pattern present in the context as well as by the global knowledge. In this case, it can be concluded that the information regarding the positions where token patterns are observed does not influence the final logit. We can easily derive the following properties.
Corollary 1.
Given a length-$T$ sequence $z_{1:T}$ where the last token is $z_{T}=q$ and the only token patterns $qv_{1}$ and $qv_{2}$ appear $f(v_{1})$ and $f(v_{2})$ times, respectively. The transformer outputs one of the following: $\operatorname*{arg\,max}_{v\in\mathcal{V}\setminus\{v_{1},v_{2}\}}\pi_{b}(v\mid q)$, $v_{1}$, or $v_{2}$.
For vocabulary $v\neq v_{1},v_{2}$, the logit $\xi_{v}$ is determined only by $\log{\pi_{b}(v\mid q)}$ and this means that the candidate for the next token is $\operatorname*{arg\,max}_{v\in\mathcal{V}\setminus\{v_{1},v_{2}\}}\pi_{b}(v\mid q)$. On the other hand, The logits for $v_{1}$ and $v_{2}$ have the other term, which makes them other candidates for the next token. Overall, Corollary 1 states that the model either uses global bigram knowledge or in-context knowledge, depending on the bigram conditionals $\pi_{b}(\cdot\mid q)$ and the frequency of token patterns $qv_{1}$ and $qv_{2}$ in the context.
5 Experiments
Henceforth, we denote by $\operatorname{TF}_{\text{ape}},\operatorname{TF}_{\text{rpe}}$ and $\operatorname{TF}_{\text{orc}}$ two-layer transformer with APE, RPE, and oracle two-layer transformer parameterized by $\tau_{1},\tau_{2}$ as in Def. 3, respectively. Note that we construct $\operatorname{TF}_{\text{orc}}$ by setting appropriate weights for each matrix in the transformer. We used the two-layer transformer architecture $\operatorname{TF}_{\text{ape}}$ constructed in Sec. 3.1.
5.1 Oversight of in-context knowledge
Figure 2: Comparison of a two-layer transformer with APE and RPE in capturing previous token information. Even when trained with sequences of length $256$, the attention block in the first layer of $\operatorname{TF}_{\text{APE}}$ fails to attend to the previous token at positions $t>128$, while $\operatorname{TF}_{\text{RPE}}$ attends to previous tokens regardless of the positions.
Training process of previous token head
We investigated whether the training of transformers directs attention towards the previous token. Specifically, $\operatorname{TF}_{\text{rpe}}$ and $\operatorname{TF}_{\text{ape}}$ were trained using inputs of length $256$ generated according to the sequence generation rule. Details regarding the hyperparameters and sequence generation process are provided in Appendix D.
For an associative memory $\sum_{(i,j)\in M}u_{i}v_{j}^{\top}$, we employed the memory recall metric [12, 18, 4]:
$R(W)=\frac{1}{\|M\|}\sum_{(i,j)\in M}\mathbf{1}\{\operatorname*{arg\,max}_{i^{\prime}}u_{i^{\prime}}^{\top}Wv_{j}=i\}.$
For example, $\operatorname{TF}_{\text{rpe}}$ calculates the attention with $W_{K}^{1}=\sum_{k\in Q}w_{E}(k)r_{-1}^{\top}$, and $W_{Q}^{1}=I$ in the first layer attention block. We want each vocabulary $w_{E}(k)$ to be paired with the information of the previous position $r_{-1}$, not other positions $r_{-i}$ where $i\neq 1$. Hence, $R(W_{K}^{1})$ is computed as
$\frac{1}{\|Q\|}\sum_{k\in Q}\mathbf{1}\{\operatorname*{arg\,max}_{i^{\prime}}r_{-i^{\prime}}^{\top}W_{K}^{1}w_{E}(k)=-1\}.$
For $\operatorname{TF}_{\text{ape}}$, we similarly have
$\frac{1}{T}\sum_{t=2}^{T}\mathbf{1}\{\operatorname*{arg\,max}_{t^{\prime}}p_{t^{\prime}}^{\top}W_{K}^{1}p_{t}=t-1\},$
where $T$ is the maximum input length for $\operatorname{TF}_{\text{ape}}$.
As we can see from Fig. 2, over $90\%$ of desired associations are stored in $W_{K}^{1}$ for $\operatorname{TF}_{\text{rpe}}$, while $\operatorname{TF}_{\text{ape}}$ fails to remember most of the associations at positions $256128<t<$ 256 even though the model was trained with sequences of length $256$.
Discussion
One of the worst-case scenarios for $\operatorname{TF}_{\text{ape}}$ occurs when a trigger token does not appear by chance in the range $t\leq 128$ and only appears for the first time at $t>128$. Although such a sequence is contrived, its probability of occurrence is non-zero. For such sequences, models with an incomplete previous token head fail to associate the trigger token’s information with the output token, resulting in incorrect predictions without leveraging in-context information. In contrast, $\operatorname{TF}_{\text{rpe}}$, which links words to their relative positions, ensures that the in-context information is not overlooked, regardless of where the trigger token first appears, thereby enabling correct predictions.
5.2 Global knowledge v.s. in-context knowledge
Figure 3: Ratio of predicting $B_{1}$ and $B_{2}$ given the prompt "$A\,B_{1}\,,\dots,\,A\,B_{1}\,,\,A\,B_{2}\,,\dots,\,A\,B_{2}\,,\,A$". The transformer shows an increasing tendency to complete the context with $B_{1}$ as the frequency of token pattern $A\,B_{1}$ in the context grows.
To validate the results derived from the theory, we conducted experiments using real data as well as data generated by the bigram model.
Analogical reasoning task
We utilize the analogical reasoning task, one of the most practical tasks for bigram analysis. Specifically, we focus on the "capital-world" analogy type of questions from the Google Analogy Dataset [36], which contains a total of 19,544 questions. This subset includes 232 unique vocabulary items. Denoting a word $A$ (e.g., Tokyo) and its corresponding analogy counterpart $A^{*}$ (e.g., Japan), the prompts were constructed in the form $A\,A^{*}\,,\,B\,B^{*}\,,\,C\,C^{*}\,,\dots$, with pairs separated by commas. Using these generated prompts, we trained $\operatorname{TF}_{\text{rpe}}$. For details on the training and dataset setup, please refer to Appendix D.
Collision of context information
As shown in Corollary 1, the model predicts either a token based on global knowledge or the token that appears in the context. To confirm this, we randomly sampled $A$ from the capitals in the Google Analogy Dataset and constructed prompts using word pairs $A\,B_{1}$ and $A\,B_{2}$ that were not learned as global knowledge. The resulting prompts were of the form $A\,B_{1}\,,\dots,\,A\,B_{1}\,,\,A\,B_{2}\,,\dots,\,A\,B_{2}\,,\,A$, which were then used for predicting the next token. We generated 1,000 prompts and calculated the proportion of cases where either $B_{1}$ or $B_{2}$ was predicted, while controlling the frequency of $B_{1}$ and $B_{2}$ in the context.
The result is shown in Fig. 3. As the number of token pattern $A\,B_{1}$ increases, the model predicts $B_{1}$ as the next token more frequently. Furthermore, it can be observed that the prediction trend reverses when the frequency of $B_{1}$ and $B_{2}$ in the context is nearly equal. Additionally, we observe from global ratio that the model predicted some vocabulary $B$ from global knowledge. Especially, the figure illustrates that it is more likely to output global knowledge when the number of $B_{1}$ and $B_{2}$ are the same, or when the number of $B_{1}$ or $B_{2}$ is almost maximum. Please refer to Appendix E.2 for further discussion on these phenomena.
6 Conclusion
We analyzed the influence of interaction between global bigram and in-context knowledge to the two-layer transformer through the lens of associative memory. We theoretically and empirically verified that relative positional encoding enables transformers to generalize longer sequences, and that how the next token prediction is conducted within the transformer that was trained to store global knowledge and to have induction head.
7 Limitations
Our study has certain limitations that should be acknowledged. First, the analysis presented in this work is conducted using a two-layer transformer model. While this allows for a controlled and interpretable exploration of the underlying mechanisms, the findings may not directly generalize to LLMs, which typically involve significantly more layers and complex interactions.
Second, while our focus on induction heads provides valuable insights into in-context learning ICL, it is important to note that ICL leverages additional computational circuits beyond induction heads. As such, our explanation captures only part of the broader mechanisms underlying ICL, leaving room for further investigation into other contributing factors.
8 Ethics statement
This work is purely theoretical, supported by controlled experiments designed to validate the proposed analyses. No personal or sensitive data were involved at any stage of this research. All experiments were conducted using datasets that are publicly available, ensuring compliance with ethical standards for data usage.
Appendix A Additional related work
In-context learning and induction head
Bayesian inference is also a popular perspective for comprehending ICL. Xie et al. [57] and Wies et al. [55] investigated how a pretrained transformer implicitly discovers the underlying concept or latent task of a prompt by leveraging a mixture of distributions. An alternative line of work [60] relates the Bayesian model averaging to the attention mechanism. Wang et al. [54] studies how information is aggregated to label words in ICL. In terms of two-layer transformer, Chen et al. [9] studied ICL on $n$-gram Markov chain, proving that gradient flow optimization under the cross-entropy loss leads to a model that exhibits a generalized induction head mechanism. Edelman et al. [15] showed two-layer attention-only transformers learn induction heads that estimate the next token’s conditional probability based on prior occurrences, enabling near Bayes-optimal performance. Similarly, Nichani et al. [39] showed that two-layer transformers learn causal structure via gradient descent, suggesting that induction head is a special case of this structure. Ren et al. [45] discovered a more generalized version of induction heads, which are referred to as semantic induction heads.
Length generalization and positional encoding
Although we point out the oversight of in-context information by a transformer with APE, this can be considered as the failure of length generalization. Since the introduction of the Transformer architecture [50], there has been a wealth of research on positional encoding [61]. Originally, Vaswani [50] adopted sinusoidal positional encoding that transforms the tokens’ absolute positions. Subsequently, Kiyono et al. [29] and Likhomanenko et al. [32] proposed shifting absolute sinusoidal positional encoding during training to achieve shift invariance. Wang et al. [52] extends word embedding as continuous functions with respect to token position, leading to smooth word representations. Meanwhile, it is demonstrated that RPE is superior to APE for longer sequence generalization in various tasks [38, 32, 23, 24]. Additionally, Sinha et al. [48] questioned whether models can learn the relative distance between words when equipped with APE. The first RPE was formulated by Shaw et al. [47], where a trainable encoding is added to the key before computing the dot product with the query. The transformer in our work also adopts this type of positional encoding, while we fix the positional encodings to randomly initialized vectors. A variety of different RPE methods were presented [23, 28] until now, but DeBERTa [19], RoPE [49], T5-Bias [44], and ALiBi [41] are the most popular choices. Ruoss et al. [46] suggested that randomizing absolute and relative positional encoding enhance the model’s ability to generalize to input sequences of lengths not encountered during training, but they used the encoder-ony Transformer, which is different from ours.
Appendix B Learning of ideal transformer
B.1 Notation table
We prepare tab. 2, which provides a concise summary of the notations used throughout the appendices.
Symbol Description
$d$ Dimensionality of the positional encoding
$\mathcal{V}$ Vocabulary set
$V$ Vocabulary size
$w_{E}(z_{t})$ Embedding of token $z_{t}$
$r_{s-t}$ Relative positional encoding expressing the relation between $s$-th token and current one
$\sigma$ softmax function
$\Phi_{1}$ product of two randomly initialized matrices $W_{O}^{1}$ and $W_{V}^{1}$
$t_{q}$ Index of the first trigger token
$t_{o}$ Index of the output token corresponding to the first trigger token, $t_{o}=t_{q}+1$
$T$ Index of the second trigger token, or the input length
$[N]$ a set of natural number up to $N$, $\{1,2,\dots,N\}$
$l$ cross entropy loss function
$\eta$ learning rate
$\tau$ constant, defined by $\mathbb{E}_{t_{o}}[\sum_{t=t_{o}}^{T}1/t]$
$\alpha$ constant, $\alpha=\eta/VT$
$\alpha^{\prime}$ constant, $\alpha=\eta\tau/VT$
$q$ a variable representing the trigger token
$\pi_{u}$ a uniform distribution from which $z_{1}$ is sampled.
$\pi_{q}$ a uniform distribution by which the trigger token is determined.
$\pi_{o}$ a uniform distribution by which the output token is determined.
$\pi_{b}$ a uniform distribution from which $z_{2:T}$ is sampled conditioned on the previous token.
Table 2: Table of Notations
B.2 Setup
Here we consider a simplified setting to analyze the influence of relative positional encoding in the training dynamics of our two-layer attention only transformer. Specifically,
  1. 1.
    Input Sequence We consider an input sequence $z_{1:T}\in\mathcal{V}^{T}$ which has one trigger token $q$ appearing twice, and ends with the trigger token. In other words, let $t_{q}$ be the first occurrence position. Then, we have $z_{t_{q}}=z_{T}=q$. From the bigram generation rule in Sec.3.2, $z_{t_{q}+1}$ and $z_{T+1}$ are the output token.
  2. 2.
    Probability Distribution Assumptions: The bigram is generated by uniform distributions over $[V]$ for any index $i$, i.e., $\pi_{u}$, $\pi_{q}$, $\pi_{o}$, and $\pi_{b}(\cdot\mid i)$ are uniformly distributed.
  3. 3.
    Simplification of Loss Function: Consider the loss only for sequences of length $T$ where the final input token $z_{T}$ is the second occurrence of the trigger token, and the label $y$ is $z_{T+1}$, the corresponding output token.
  4. 4.
    Simplification for Learning Focus: Our approach involves sequentially training $W_{2}^{O}$, $W_{2}^{K}$, and $W_{1}^{K}$ from top to bottom. We employ zero-initialization and carry out a single gradient descent step.
  5. 5.
    Initialization and Freezing: To achieve our goal of showing that $W_{O}^{2},W_{K}^{2}$ and $W_{K}^{1}$ learn to be an associative memory, we zero-initialize the three matrices. For other matrices such as $W_{V}^{2},W_{V}^{1}$ and $W_{O}^{1}$ are randomly initialized from Gaussian distribution. We set $W_{Q}^{1}$ and $W_{Q}^{2}$ to identity matrix. All these matrices except for $W_{O}^{2},W_{K}^{2}$ and $W_{K}^{1}$ are freezed during the training process.
B.3 Theoretical analysis of learning
Theorem 3 (formal).
Under the setup described in Sec. B.2, a two-layer attention only transformer with relative positional encoding learns associations of the associative memory transformer, i.e., the memories $W_{O}^{2},W_{K}^{2}$ and $W_{K}^{1}$ can be written as the sum of the terms that constitute the associative memory transformer.
$W_{K}^{1}$
$=\sum_{k\in Q}\chi(k)w_{E}(k)r_{-1}^{\top}+(W_{K}^{1})^{\prime},$ (14)
$W_{K}^{2}$
$=\sum_{k\in Q}\psi(k)w_{E}(k)(W_{O}^{1}W_{V}^{1}w_{E}(k))^{\top}+(W_{K}^{2})^{\prime},$ (15)
$W_{O}^{2}$
$=\sum_{v=1}^{V}\omega(v)w_{U}(v)(W_{V}^{2}w_{E}(v))^{\top}+(W_{O}^{2})^{\prime},$ (16)
where $\chi,\psi:\mathcal{V}\to\mathbb{R}$ and $\omega:\mathbb{N}\to\mathbb{R}$ denotes the learned score of each association, and $(W_{K}^{1})^{\prime},(W_{K}^{2})^{\prime}$ and $(W_{O}^{2})^{\prime}$ are the other associations.
Proof.
We will demonstrate that associative memory can be learned by performing one gradient descent step in a top-down manner, starting with $W_{O}^{2}$, followed by $W_{K}^{2}$, and then $W_{K}^{1}$.
Step1: Training of $W_{O}^{2}$
We begin with the dynamics of $W_{O}^{2}$. The input to the second layer attention, $x^{(1)}_{t}$, is composed of the sum of the values from the first layer attention, $x^{(1,0)}_{t}$, and the values from the residual connection, $x^{(1,1)}_{t}$. Since $W_{K}^{1}$ is initialized to zero, the attention of the first layer is evenly distributed across all the tokens $z_{1:t}$ and we have
$x^{(1,0)}_{t}=\frac{1}{t}\sum_{s=1}^{t}\Phi_{1}w_{E}(z_{s}),$
where $\Phi_{1}=W_{O}^{1}W_{V}^{1}$. Additionally, the residual stream carries the token embedding:
$x^{(1,1)}_{t}=w_{E}(z_{t}).$
Since the matrix $W_{K}^{2}$ is also zero-initialized, the value of the second layer attention can be computed through
$x^{(2)}_{t}=W_{V}^{2}\left(\frac{1}{t}\sum_{s=1}^{t}x^{(1)}_{s}\right).$
Now the logit prediction is given by $W_{U}(W_{O}^{2}x^{(2)}_{T}+x^{(1)}_{T})$ with the residual connection. We note that when $d$ is large, thanks to the near-orthogonality, $W_{U}x_{T}^{(1)}=W_{U}\left(\sum_{s=1}^{T}\Phi_{1}w_{E}(z_{s})/T+w_{E}(z_{T})\right)$ is negligible because $x_{T}^{(1)}$ does not contain $w_{U}(v)$ for any $v\in\mathcal{V}$. This enables us to use Lemma 2 and one gradient step with learning-rate $\eta$ yields
$W_{O}^{2}$
$=\frac{\eta}{V}\sum_{k=1}^{V}w_{U}(k)\mathbb{E}[x^{(2)}_{T}\mid y=k]^{\top}$
$\quad-\frac{\eta}{V}\sum_{k=1}^{V}w_{U}(k)\mathbb{E}\left[\frac{\hat{p}_{W}(k\mid x)}{p(y=k)}x^{(2)}_{T}\right]^{\top}$
$=\frac{\eta}{V}\sum_{k=1}^{V}w_{U}(k)(\mathbb{E}[x^{(2)}_{T}\mid y=k]-\mathbb{E}[x^{(2)}_{T}])^{\top}.$ (17)
This transformation comes from the assumption that every token is sampled from a uniform distribution ($p(y=k)=1/V)$ and all the logits are 00 when $W_{O}^{2}=O_{d\times d}$, which means that $\hat{p}_{W}(k\mid x)=1/V$.
Given that $y=k$, it holds that the first output token $z_{t_{o}}=k$. For $z_{i}$ ($i\neq t_{o}$), the uniform distribution can possibly generate any token, independently of the condition $y=k$. Since we have
$x_{T}^{(2)}=\frac{1}{T}\sum_{t=1}^{T}W_{V}^{2}\left(w_{E}(z_{t})+\frac{1}{t}\sum_{s=1}^{t}\Phi_{1}w_{E}(z_{s})\right),$
the expression $\mathbb{E}\left[x^{(2)}_{T}\mid y=k\right]-\mathbb{E}\left[x^{(2)}_{T}\right]$ leaves only the terms related to $z_{t_{o}}$. This yields
$\mathbb{E}\left[x^{(2)}_{T}\mid y=k\right]-\mathbb{E}\left[x^{(2)}_{T}\right]$
$=\frac{1}{T}W_{V}^{2}(\mathbb{E}[w_{E}(z_{t_{o}})\mid y=k]-\mathbb{E}[w_{E}(z_{t_{o}})])$
$\quad+\frac{1}{T}\cdot\mathbb{E}\left[\sum_{t=t_{o}}^{T}W_{V}^{2}\frac{1}{t}\Phi_{1}w_{E}(z_{t_{o}})\mid y=k\right]$
$\quad-\frac{1}{T}\cdot\mathbb{E}\left[\sum_{t=t_{o}}^{T}W_{V}^{2}\frac{1}{t}\Phi_{1}w_{E}(z_{t_{o}})\right]$ (18)
$=\frac{1}{T}W_{V}^{2}(w_{E}(k)-\mathbb{E}[w_{E}(z_{t_{o}})])$
$\quad+\frac{\tau}{T}W_{V}^{2}\Phi_{1}(w_{E}(k)-\mathbb{E}[w_{E}(z_{t_{o}})]),$ (19)
where $\tau:=\mathbb{E}\left[\sum_{t=t_{o}}^{T}\frac{1}{t}\right]$. Since each token is generated from a uniform distribution, the expected value of $w_{E}(z_{t_{o}})$, conditioned on nothing, is given by:
$\mathbb{E}\left[w_{E}(z_{t_{o}})\right]=\frac{1}{V}\sum_{k=1}^{V}w_{E}(k).$
Combining eq. 17 and B.3, the near-orthogonality of Gaussian vectors gives us
$w_{U}(k)^{\top}W_{O}^{2}W_{V}^{2}w_{E}(j)$
$\approx\frac{\eta}{VT}\mathbf{1}(k=j)+O\left(\frac{\eta}{V^{2}T}\right),$ (20)
$w_{U}(k)^{\top}W_{O}^{2}W_{V}^{2}\Phi_{1}w_{E}(j)$
$\approx\frac{\eta\tau}{VT}\mathbf{1}(k=j)+O\left(\frac{\eta\tau}{V^{2}T}\right).$ (21)
This is the same result as in the proof from Bietti et al. [4]. So far, the weight matrix $W_{O}^{2}$ is trained to perform as an associative memory so that
$w_{U}(k)^{\top}W_{O}^{2}W_{V}^{2}w_{E}(z_{t})\approx\alpha\mathbf{1}\{z_{t}=k\},$ (22)
with a constant $\alpha=\frac{\eta}{TV}$.
Step2: Training of $W_{K}^{2}$
Since the training of $W_{O}^{2}$ is finished, we have Eq. B.3. When $W_{K}^{1}=W_{K}^{2}=O_{d\times d}$, Lemma 4 demonstrates that the next token prediction is distributed among all tokens, which means that $\hat{p}(v\mid x)=1/V$ for any vocabulary $v$. According to Lemma 5, one step of gradient descent yields
$W_{K}^{2}$
$=\frac{\eta}{TN}\sum_{k,t}\mathbb{E}\Bigl{[}w_{U}(k)^{\top}\Phi_{2}x_{t}^{(1)}\cdot(x_{t}^{(1)}-\bar{x}^{(1)})$
×$\qquad\times(x_{T}^{(1)})^{\top}\mid y=k\Bigr{]}$
$\quad-\frac{\eta}{TN}\sum_{k,t}\mathbb{E}\Bigl{[}w_{U}(k)^{\top}\Phi_{2}x_{t}^{(1)}\cdot(x_{t}^{(1)}-\bar{x}^{(1)})$
×$\qquad\times(x_{T}^{(1)})^{\top}\Bigr{]},$ (23)
with $\bar{x}^{(1)}=\frac{1}{T}\sum_{t=1}^{T}x_{t}^{(0)}$.
We simplify our transformer architecture by setting $x_{t}^{(1,1)}$ as the queries and values, and $x_{t}^{(1,0)}$ as the keys. Furthermore, we add the condition that the trigger token $z_{t_{q}}=j$ and rewrite eq. 23. This conditioning leverages the fact that each token in $z_{t_{q}}$ is sampled from a uniform distribution. Now we have
$W_{K}^{2}$
$=\frac{\eta}{TN}\sum_{k,t}\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}\Bigl{[}w_{U}(k)^{\top}\Phi_{2}x_{t}^{(1,1)}$
×$\qquad\times(x_{t}^{(1,0)}-\bar{x}^{(1,0)})(x_{T}^{(1,1)})^{\top}\mid y=k,z_{t_{q}}=j\Bigr{]}$
$\quad-\frac{\eta}{TN}\sum_{k,t}\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}\Bigl{[}w_{U}(k)^{\top}\Phi_{2}x_{t}^{(1,1)}$
×$\qquad\times(x_{t}^{(1,0)}-\bar{x}^{(1,0)})(x_{T}^{(1,1)})^{\top}\mid z_{t_{q}}=j\Bigr{]},$
where $\bar{x}^{(1,0)}=\frac{1}{T}\sum_{t=1}^{T}x_{t}^{(1,0)}$. Using the formula 22, we obtain
$W_{K}^{2}$
$\approx\frac{\alpha\eta}{TN^{2}}\sum_{j=1}^{N}\sum_{k=1}^{N}\mathbb{E}\Biggl{[}\sum_{t=1}^{T}\mathbbm{1}\{z_{t}=k\}$
×$\qquad\times(x_{t}^{(1,0)}-\bar{x}^{(1,0)})w_{E}(z_{T})^{\top}\mid y=k,z_{t_{q}}=j\Biggr{]}$
$\quad-\frac{\alpha\eta}{TN^{2}}\sum_{j=1}^{N}\sum_{k=1}^{N}\mathbb{E}\Biggl{[}\sum_{t=1}^{T}\mathbbm{1}\{z_{t}=k\}$
×$\qquad\times(x_{t}^{(1,0)}-\bar{x}^{(1,0)})w_{E}(z_{T})^{\top}\mid z_{t_{q}}=j\Biggr{]}$
$=\frac{\alpha\eta}{TN^{2}}\sum_{j=1}^{N}\sum_{k=1}^{N}\Delta_{k,j}w_{E}(j)^{\top},$ (24)
where $\Delta_{k,j}$ is defined by
$\Delta_{k,j}$
$:=\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{z_{t}=k\}\right.$
×$\quad\left.\times(x_{t}^{(1,0)}-\bar{x}^{(1,0)})\mid y=k,z_{t_{q}}=j\right]$
$\quad-\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{z_{t}=k\}\right.$
×$\quad\left.\times(x_{t}^{(1,0)}-\bar{x}^{(1,0)})\mid z_{t_{q}}=j\right].$
The sum in $\Delta_{k,j}$ can be partitioned into three distinct groups: (i) $\Delta_{k,j}^{o}$, where $t$ is the index of the output token; (ii) $\Delta_{k,j}^{q}$, where $t$ corresponds to the indices of the trigger tokens; and (iii) $\Delta_{k,j}^{r}$, which includes all other cases. Mathematically, let $t_{o}\geq 2$ be a random variable, and $t_{q}=t_{o}-1$, we write
$\Delta_{k,j}^{o}$
$:=\mathbb{E}\left[\mathbbm{1}\{z_{t_{o}}=k\}(x_{t_{o}}^{(1,0)}-\bar{x}^{(1,0)})\mid y=k,z_{t_{q}}=j\right]$
$\quad-\mathbb{E}\left[\mathbbm{1}\{z_{t_{o}}=k\}(x_{t_{o}}^{(1,0)}-\bar{x}^{(1,0)})\mid z_{t_{q}}=j\right],$
$\Delta_{k,j}^{q}$
$:=\mathbb{E}\left[\sum_{t\in\mathcal{T}_{q}}\mathbbm{1}\{z_{t}=k\}(x_{t}^{(1,0)}-\bar{x}^{(1,0)})\mid y=k,z_{t_{q}}=j\right]$
$\quad-\mathbb{E}\left[\sum_{t\in\mathcal{T}_{q}}\mathbbm{1}\{z_{t}=k\}(x_{t}^{(1,0)}-\bar{x}^{(1,0)})\mid z_{t_{q}}=j\right],$
$\Delta_{k,j}^{r}$
$:=\mathbb{E}\left[\sum_{t\in\mathcal{T}_{r}}\mathbbm{1}\{z_{t}=k\}(x_{t}^{(1,0)}-\bar{x}^{(1,0)})\right.$
$\qquad\qquad\left.\mid y=k,z_{t_{q}}=j\right]$
$\quad-\mathbb{E}\left[\sum_{t\in\mathcal{T}_{r}}\mathbbm{1}\{z_{t}=k\}(x_{t}^{(1,0)}-\bar{x}^{(1,0)})\mid z_{t_{q}}=j\right],$
where $\mathcal{T}_{q}=\{t_{q},T\}$ and $\mathcal{T}_{r}=\{1\leq i\leq T\mid i\in\mathbb{N},i\neq t_{o},i\neq t_{q},i\neq T\}$.
First, we will take a look at $\Delta_{k,j}^{o}$. Note that $\mathbb{E}[\mathbbm{1}\{z_{t_{o}}=k\}\mid y=k,z_{t_{q}}=j]=1$ and $\mathbb{E}[\mathbbm{1}\{z_{t_{o}}=k\}\mid z_{t_{q}}=j]=1/N$, so we find that
$\Delta_{k,j}^{o}$
$=\left(1-\frac{1}{N}\right)\mathbb{E}\left[x_{t}^{(1,0)}-\bar{x}^{(1,0)}\mid y=k\right]$
$=\frac{N-1}{N}\mathbb{E}\left[\frac{1}{t_{o}}\sum_{s=1}^{t_{o}}\Phi_{1}w_{E}(z_{s})\right.$
$\qquad\left.-\frac{1}{T}\sum_{t=1}^{T}\frac{1}{t}\sum_{s=1}^{t}\Phi_{1}w_{E}(z_{s})\right]$
$=\frac{N-1}{N}\sum_{i=1}^{N}a_{k,j,i}\Phi_{1}w_{E}(i),$ (25)
where $a_{k,j,i}$ is the coefficient of $\Phi_{1}w_{E}(i)$ by calculating
$\mathbb{E}\left[\frac{1}{t_{o}}\sum_{s=1}^{t_{o}}\Phi_{1}w_{E}(z_{s})-\frac{1}{T}\sum_{t=1}^{T}\frac{1}{t}\sum_{s=1}^{t}\Phi_{1}w_{E}(z_{s})\right].$
The analysis of $a_{k,j,i}$ is done by the work [4]:
$\frac{1}{N}\sum_{k=1}^{N}a_{k,j,i}\approx\begin{cases}O\left(\frac{1}{N}\right),&\text{if }i\neq j,\\ \Omega\left(\frac{1}{T}\right),&\text{otherwise}.\end{cases}$
Thus, we have the following approximation:
$(\Phi_{1}w_{E}(i))^{\top}\left(\frac{1}{N}\sum_{k=1}^{N}\Delta_{k,j}^{o}\right)$
$\approx\begin{cases}O\left(\frac{1}{N}\right),&\text{if }i\neq j,\\ \Omega\left(\frac{1}{T}\right),&\text{if }i=j.\end{cases}$ (26)
Similarly, we have $\Delta_{k,j}^{q}=O\left(\frac{1}{N}\right)$ and $\Delta_{k,j}^{r}=O\left(\frac{T}{N}\right)$ based on the discussion in Bietti et al. [4]. By combining these analysis with Eq. B.3, we multiply $(\Phi_{1}w_{E}(i))^{\top}$ from left and $w_{E}(j)$ from right to Eq. 24 and we establish that
$(\Phi_{1}w_{E}(i))^{\top}W_{K}^{2}w_{E}(j)$
$\approx\frac{\alpha\eta}{TN}\left\{\Omega\left(\frac{1}{T}\right)\mathbbm{1}\{i=j\}+O\left(\frac{T}{N}\right)\right\}$ (27)
Step3: Training of $W_{K}^{1}$
So far, the gradient descent step as to $W_{O}^{2}$ and $W_{K}^{2}$ gives them the ability to behave as an associative memory:
$w_{U}(k)^{\top}W_{O}^{2}W_{V}^{2}w_{E}(z_{t})$ $=\alpha\mathbf{1}\{z_{t}=k\},$ (28)
$(\Phi_{1}w_{E}(z_{t}))^{\top}W_{K}^{2}w_{E}(z_{T})$ $=\alpha^{\prime}\mathbf{1}\{z_{t}=z_{T}\}.$ (29)
From Lemma 3, we admit that the model predicts the next token almost randomly, i.e., $\hat{p}(v\mid x)=1/V$ for any vocabulary $v$, when $W_{K}^{0}=O_{d\times d}$. This enables us to employ Lemma 5 and $W_{K}^{1}$ after one gradient descent step gives the following:
$W_{K}^{1}$
$=\frac{\eta}{V}\sum_{k=1}^{N}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}w_{U}(k)^{\top}\Phi_{2}w_{E}(z_{t})\right.$
×$\quad\times\frac{1}{t}\sum_{s=1}^{t}(\Phi_{1}w_{E}(z_{s}))^{\top}W_{K}^{2}$
×$\quad\left.\times w_{E}(z_{T})q_{s,t}w_{E}(z_{t})^{\top}\mid y=k\right]$
$\quad-\frac{\eta}{V}\sum_{i=k}^{N}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}w_{U}(k)^{\top}\Phi_{2}w_{E}(z_{t})\right.$
×$\quad\left.\times\frac{1}{t}\sum_{s=1}^{t}(\Phi_{1}w_{E}(z_{s}))^{\top}W_{K}^{2}w_{E}(z_{T})q_{s,t}w_{E}(z_{t})^{\top}\right]$
$\quad-\frac{\eta}{V}\sum_{k=1}^{N}\mathbb{E}\left[\left(w_{U}(k)^{\top}\Phi_{2}\bar{x}_{1:T}\right)\right.$
×$\quad\times\frac{1}{T}\sum_{t=1}^{T}\frac{1}{t}\sum_{s=1}^{t}(\Phi_{1}w_{E}(z_{s}))^{\top}W_{K}^{2}w_{E}(z_{T})$
×$\quad\left.\times q_{s,t}w_{E}(z_{t})^{\top}\mid y=k\right]$
$\quad+\frac{\eta}{V}\sum_{k=1}^{N}\mathbb{E}\left[\left(w_{U}(k)^{\top}\Phi_{2}\bar{x}_{1:T}\right)\right.$
×$\quad\times\frac{1}{T}\sum_{t=1}^{T}\frac{1}{t}\sum_{s=1}^{t}(\Phi_{1}w_{E}(z_{s}))^{\top}W_{K}^{2}w_{E}(z_{T})$
×$\quad\left.\times q_{s,t}w_{E}(z_{t})^{\top}\right],$
with $q_{s,t}=(w_{E}(z_{s})+r_{s-t})-(\bar{w}_{E}(z_{1:t})+\bar{r}_{1-t:0})$ and $\bar{w}_{E}(z_{1:t})=\frac{1}{t}\sum_{i=1}^{t}w_{E}(z_{i})$. Now, fix an arbitrary vocabulary $v\in\mathcal{V}$. Given that $w_{U}(k)^{\top}\Phi_{2}w_{E}(z_{t})=\alpha\mathbf{1}\{z_{t}=k\}$ and that $(\Phi_{1}w_{E}(z_{s}))^{\top}W_{K}^{2}w_{E}(z_{T})=\alpha^{\prime}\mathbf{1}\{z_{s}=z_{T}\}$, which equals $\alpha^{\prime}$ when $s=t_{q}$ or $s=T$, we manipulate the expression:
$W_{K}^{1}$
$=\frac{\eta}{V}\sum_{k=1}^{V}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\alpha\mathbf{1}\{z_{t}=k\}\right.$
×$\quad\times\frac{\alpha^{\prime}}{t}(\mathbf{1}\{t_{q}\leq t\}q_{t_{q},t}+\mathbf{1}\{t=T\}q_{T,T})$
×$\quad\left.\times w_{E}(z_{t})^{\top}\mid y=k\right]$
$\quad-\frac{\eta}{V}\sum_{k=1}^{V}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\alpha\mathbf{1}\{z_{t}=k\}\right.$
×$\quad\times\frac{\alpha^{\prime}}{t}(\mathbf{1}\{t_{q}\leq t\}q_{t_{q},t}+\mathbf{1}\{t=T\}q_{T,T})$
×$\quad\left.\times w_{E}(z_{t})^{\top}\right]$
$\quad-\frac{\eta}{V}\sum_{k=1}^{V}\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t}^{\prime}=k\}\right.$
×$\quad\times\frac{1}{T}\sum_{t=1}^{T}\frac{\alpha^{\prime}}{t}(\mathbf{1}\{t_{q}\leq t\}q_{t_{q},t}+\mathbf{1}\{t=T\}q_{T,T})$
×$\quad\left.\times w_{E}(z_{t})^{\top}\mid y=k\right]$
$\quad+\frac{\eta}{V}\sum_{k=1}^{V}\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t}^{\prime}=k\}\right.$
×$\quad\times\frac{1}{T}\sum_{t=1}^{T}\frac{\alpha^{\prime}}{t}(\mathbf{1}\{t_{q}\leq t\}q_{t_{q},t}+\mathbf{1}\{t=T\}q_{T,T})$
×$\quad\left.\times w_{E}(z_{t})^{\top}\right].$
With the near-orthogonality of the token embeddings, we write
$W_{K}^{1}w_{E}(v)$ (30)
$=\frac{\eta\alpha\alpha^{\prime}}{VT}\sum_{k=1}^{V}\sum_{t=1}^{T}\frac{1}{t}(A_{t,k}-B_{t,k}-C_{t,k}+D_{t,k})$
$\quad+\frac{\eta\alpha\alpha^{\prime}}{VT^{2}}\sum_{k=1}^{V}(A^{\prime}_{T,k}-B^{\prime}_{T,k}-C^{\prime}_{T,k}+D^{\prime}_{T,k}).$ (31)
where, we define each term as follows:
$A_{t,k}$ $=\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}\right.$
×$\qquad\left.\times q_{t_{q},t}\mathbf{1}\{z_{t}=v\}\mid y=k\right],$
$B_{t,k}$ $=\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}q_{t_{q},t}\mathbf{1}\{z_{t}=v\}\right],$
$C_{t,k}$ $=\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t}^{\prime}=k\}\mathbf{1}\{t_{q}\leq t\}\right.$
×$\qquad\left.\times q_{t_{q},t}\mathbf{1}\{z_{t}=v\}\mid y=k\right],$
$D_{t,k}$ $=\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t}^{\prime}=k\}\mathbf{1}\{t_{q}\leq t\}\right.$
×$\qquad\left.\times q_{t_{q},t}\mathbf{1}\{z_{t}=v\}\right],$
$A^{\prime}_{T,k}$ $=\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\mid y=k\right],$
$B^{\prime}_{T,k}$ $=\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\right],$
$C^{\prime}_{T,k}$ $=\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t}^{\prime}=k\}\right.$
×$\qquad\left.\times q_{T,T}\mathbf{1}\{z_{T}=v\}\mid y=k\right],$
$D^{\prime}_{T,k}$ $=\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t}^{\prime}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\right].$
We will further split these four terms based on relative positional encoding and token embedding that are contained in $q_{t_{q},t}$. For example, we decompose $A_{t,k}$ into $A_{t,k}^{R}$ and $A_{t,k}^{z}$ by defining
$A_{t,k}^{R}$
$=\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}\left(r_{t_{q}-t}-\bar{r}_{1-t:0}\right)\right.$
×$\qquad\left.\times\mathbf{1}\{z_{t}=v\}\mid y=k\right],$
$A_{t,k}^{z}$
$=\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}\left(w_{E}(z_{t_{q}})-\bar{w}_{E}(z_{1:t})\right)\right.$
×$\qquad\left.\times\mathbf{1}\{z_{t}=v\}\mid y=k\right].$
(i) when $v=k$
We begin with $A_{t,k}^{R}$:
$A_{t,k}^{R}$
$=\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}\left(r_{t_{q}-t}-\bar{r}_{1-t:0}\right)\right.$
×$\qquad\left.\times\mathbf{1}\{z_{t}=v\}\mid y=k\right]$
$=\sum_{s=1}^{t}P(t_{q}=s\mid y=k)$
×$\qquad\times\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s\right](r_{s-t}-\bar{r}_{1-t:0})$
$=P(t_{q}=t-1)(r_{-1}-\bar{r}_{1-t:0})$
$\quad+\mathbf{1}\{t=T\}P(t_{q}=t)(r_{0}-\bar{r}_{1-t:0})$
$\quad+\frac{\mathbf{1}\{t=T\}}{V}\sum_{s\in[t-2]\cup\{t\}}P(t_{q}=s)(r_{s-t}-\bar{r}_{1-t:0})$
$=P(t_{q}=t-1)(r_{-1}-\bar{r}_{1-t:0})$
$\quad+\mathbf{1}\{t=T\}P(t_{q}=t)(r_{0}-\bar{r}_{1-t:0})+O\left(\frac{1}{V}\right).$
Here, we used Lemma 6 and 7 for the calculation of the expectation $\mathbb{E}[\mathbf{1}\{z_{t}=k\}\mid\cdot]$.
Similarly, we will examine $B_{t,k}^{R},C_{t,k}^{R}$ and $D_{t,k}^{R}$ as well.
$B_{t,k}^{R}$
$=\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}\right.$
×$\qquad\left.\times\left(r_{t_{q}-t}-\bar{r}_{1-t:0}\right)\mathbf{1}\{z_{t}=v\}\right]$
$=\sum_{s=1}^{t}P(t_{q}=s)$
×$\qquad\times\mathbb{E}[\mathbf{1}\{z_{t}=k\}\mid t_{q}=s](r_{s-t}-\bar{r}_{1-t:0})$
$=\sum_{s=1}^{t}\frac{P(t_{q}=s)}{V}(r_{s-t}-\bar{r}_{1-t:0})$
$=O\left(\frac{1}{V}\right).$
For the term $C_{t,k}^{R}$, we make use of Lemma 8 and 9.
$C_{t,k}^{R}$
$=\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\mathbf{1}\{t_{q}\leq t\}\right.$
×$\qquad\left.\times\left(r_{t_{q}-t}-\bar{r}_{1-t:0}\right)\mathbf{1}\{z_{t}=v\}\mid y=k\right]$
$=\frac{1}{T}\sum_{s=1}^{t}P(t_{q}=s)(r_{s-t}-\bar{r}_{1-t:0})$
×$\quad\times\mathbb{E}\left[\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\right.$
×$\quad\left.\times\mathbf{1}\{z_{t}=v\}\mid y=k,t_{q}=s\right]$
$=\frac{P(t_{q}=t-1)}{T}(r_{-1}-\bar{r}_{1-t:0})$
×$\quad\times\left\{(1+\mathbf{1}\{t=T\})+\frac{(T-3)\cdot\mathbf{1}\{t\neq T\}}{V-1}\right\}$
$\quad+\frac{2P(t_{q}=t)}{T}(r_{0}-\bar{r}_{1-t:0})$
$\quad+\sum_{s=1}^{t-2}\frac{P(t_{q}=s)}{T}O\left(\frac{1}{V}\right)$
$=\frac{1}{T}\left\{P(t_{q}=t-1)\left(1+\mathbf{1}\{t=T\}\right)\right.$
×$\qquad\left.\times\left(r_{-1}-\bar{r}_{1-t:0}\right)+2P(t_{q}=t)\left(r_{0}-\bar{r}_{1-t:0}\right)\right.$
$\qquad\left.+O\left(\frac{1}{V}\right)\right\}.$
As for $D_{t,k}^{R}$, the condition $y=k$ does not exist, and $z_{t}$ follows a uniform distribution. Therefore, we obtain
$D_{t,k}^{R}$
$=\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\mathbf{1}\{t_{q}\leq t\}\right.$
×$\quad\times\left.\left(r_{t_{q}-t}-\bar{r}_{1-t:0}\right)\mathbf{1}\{z_{t}=v\}\right]$
$=\sum_{s=1}^{t}\frac{P(t_{q}=s)}{T}(r_{s-t}-\bar{r}_{1-t:0})$
×$\quad\times\mathbb{E}\left[\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\mathbf{1}\{z_{t}=v\}\mid y=k,t_{q}=s\right]$
$=O\left(\frac{1}{V}\right).$
Next, we will examine $A_{t,k}^{z},B_{t,k}^{z},C_{t,k}^{z}$ and $D_{t,k}^{z}$ in this order. Using lemma 10 and 11, we obtain.
$A_{t,k}^{z}$
$=\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}\right.$
×$\quad\left.\times(x_{t_{q}}-\bar{x}_{1:t})\mathbf{1}\{z_{t}=v\}\mid y=k\right]$
$=\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}x_{t_{q}}\mid y=k\right]$
$\quad-\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}\bar{x}_{1:t}\mid y=k\right]$
$=P(t_{q}=t-1)\mathbf{1}\{t=T\}w_{E}(k)$
$\quad+P(t_{q}=t)\mathbf{1}\{t=T-1\})w_{E}(k)$
$\quad-\frac{P(t_{q}=t-1)}{t}(1+\mathbf{1}\{t=T\})w_{E}(k)$
$\quad-\frac{\mathbf{1}\{t=T-1\})P(t_{q}=t)}{t}w_{E}(k)+O\left(\frac{1}{V}\right).$
We also get
$B_{t,k}^{z}$
$=\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}(x_{t_{q}}-\bar{x}_{1:t})\mathbf{1}\{z_{t}=v\}\right]$
$=\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}x_{t_{q}}\right]$
$\quad-\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}\bar{x}_{1:t}\right]$
$=P(z_{t}=k)\mathbb{E}[\mathbf{1}\{t_{q}\leq t\}x_{t_{q}}\mid z_{t}=k]$
$\quad-P(z_{t}=k)\mathbb{E}\left[\mathbf{1}\{t_{q}\leq t\}\bar{x}_{1:t}\mid z_{t}=k\right]$
$=O\left(\frac{1}{V}\right).$
Now we move to $C_{t,k}^{z}$. From lemma 12 and 13, it follows that
$C_{t,k}^{z}$
$=\mathbb{E}\left[\frac{1}{T}\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\mathbf{1}\{t_{q}\leq t\}\right.$
×$\qquad\times\left.(x_{t_{q}}-\bar{x}_{1:t})\mathbf{1}\{z_{t}=v\}\mid y=k\right]$
$=\mathbb{E}\left[\frac{1}{T}\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\right.$
×$\qquad\left.\times\mathbf{1}\{t_{q}\leq t\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k\right]$
$\quad-\mathbb{E}\left[\frac{1}{T}\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\right.$
×$\qquad\left.\times\mathbf{1}\{t_{q}\leq t\}\bar{x}_{1:t}\mathbf{1}\{z_{t}=k\}\mid y=k\right]$
$=w_{E}(k)\left\{\frac{2P(t_{q}=t-1)}{T}\mathbf{1}\{t=T\}\right.$
$\quad+\frac{2P(t_{q}=t)}{T}\mathbf{1}\{t=T-1\}$
$\quad-\frac{P(t_{q}=t-1)}{T}\frac{4\cdot\mathbf{1}\{t=T\}+\mathbf{1}\{t\neq T\}}{t}$
$\quad\left.-\frac{P(t_{q}=t)}{T}\frac{2\cdot\mathbf{1}\{t=T-1\}}{t}\right\}$
$\quad+\frac{1}{T}\cdot O\left(\frac{1}{V}\right).$
For $D_{t,k}^{z}$, we have
$D_{t,k}^{z}$
$=\mathbb{E}\left[\frac{1}{T}\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\right.$
×$\quad\left.\times\mathbf{1}\{t_{q}\leq t\}(x_{t_{q}}-\bar{x}_{1:t})\mathbf{1}\{z_{t}=v\}\right]$
$=P(z_{t}=v)\mathbb{E}\left[\frac{1}{T}\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\right.$
×$\quad\left.\times\mathbf{1}\{t_{q}\leq t\}(x_{t_{q}}-\bar{x}_{1:t})\mid z_{t}=v\right]$
$=O\left(\frac{1}{V}\right),$
because of the term $\mathbf{1}\{z_{t}=v\}$ under no condition. So far, we have calculated $A_{t,k},B_{t,k},C_{t,k}$ and $D_{t,k}$ for $v=k$. Now, we turn our attention to $v\neq k$.
(ii) when $v\neq k$
In this case, it is guaranteed that $\mathbf{1}\{z_{t}=k\}\neq\mathbf{1}\{z_{t}=v\}$, and thus, we have $A_{t,k}=B_{t,k}=0$. We only need to examine $C_{t,k}$ and $D_{t,k}$. We use the fact that the token $z_{t}$ has to be sampled uniformly from a potentially reduced vocabulary space with some words excluded depending on the position of $t$, we obtain
$C_{t,k}$
$=\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t}^{\prime}=k\}\mathbf{1}\{t_{q}\leq t\}\right.$
×$\qquad\left.\times q_{t_{q},t}\mathbf{1}\{z_{t}=v\}\mid y=k\right]$
$=P(z_{t}=v\mid y=k)$
×$\quad\times\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t}^{\prime}=k\}\mathbf{1}\{t_{q}\leq t\}\right.$
×$\qquad\left.\times q_{t_{q},t}\mid y=k,z_{t}=v\right]$
$=O\left(\frac{1}{V}\right),$
and
$D_{t,k}$
$=\mathbb{E}\left[\frac{1}{T}\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\right.$
×$\quad\left.\times\mathbf{1}\{t_{q}\leq t\}(x_{t_{q}}-\bar{x}_{1:t})\mathbf{1}\{z_{t}=v\}\right]$
$=P(z_{t}=v\mid y=k)$
×$\quad\times\mathbb{E}\left[\frac{1}{T}\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\mathbf{1}\{t_{q}\leq t\}\right.$
×$\qquad\left.\times q_{t_{q},t}\mathbf{1}\{z_{t}=v\}\right]$
$=O\left(\frac{1}{V}\right).$
So far, we have seen the content of $A_{t,k},B_{t,k},C_{t,k}$ and $D_{t,k}$. Now, we will shift our focus to $A^{\prime}_{T,k},B^{\prime}_{T,k},C^{\prime}_{T,k}$ and $D^{\prime}_{T,k}$. We recall that
$q_{s,t}=(w_{E}(z_{s})+r_{s-t})-(\bar{w}_{E}(z_{1:t})+\bar{r}_{1-t:0}).$
By substituting $s=T$ and $t=T$, we have
$q_{T,T}=(w_{E}(z_{T})+r_{0})-(\bar{w}_{E}(z_{1:T})+\bar{r}_{1-T:0}).$
Also, $\mathbf{1}\{z_{T}=v\}$ implies that the trigger token is $v$. we use Lemma 14 and obtain the following:
$A^{\prime}_{T,k}$
$=\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\mid y=k\right]$
=$=\begin{cases}P(t_{q}=T-1)\left(r_{0}-\bar{r}_{1-T:0}\right)\\ \quad+P(t_{q}=T-1)\frac{T-2}{T}w_{E}(k)\\ \quad+O\left(\frac{1}{V}\right)&\text{if $v=k$},\\ 0&\text{otherwise}.\end{cases}$
Similarly to the argument of $B_{t,k}$ and $D_{t,k}$, the trigger token is chosen uniformly, and thus, we have $P(z_{T}=v)=\frac{1}{V}$. This gives
$B^{\prime}_{T,k}$
$=\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\right]$
=$=\begin{cases}O\left(\frac{1}{V}\right)&\text{if $v=k$},\\ 0&\text{otherwise},\end{cases}$
and
$D^{\prime}_{T,k}$
$=\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\right]$
=$=\begin{cases}O\left(\frac{1}{V}\right)&\text{if $v=k$},\\ 0&\text{otherwise},\end{cases}$
Finally, according to Lemma 15, it is established that
$C^{\prime}_{T,k}$
$=\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\mid y=k\right]$
=$=\begin{cases}\frac{2P(t_{q}=T-1)}{T}\left(r_{0}-\bar{r}_{1-T:0}\right)\\ \quad+\frac{2P(t_{q}=T-1)}{T}\frac{T-2}{T}w_{E}(k)\\ \quad+\frac{1}{T}\cdot O\left(\frac{1}{V}\right)&\text{if $v=k$},\\ \frac{1}{T}O\left(\frac{1}{V}\right)&\text{otherwise}.\end{cases}$
To sum up, we get the following associative memory behaviors by examining Eq. 30. We point out that $\bar{r}_{1-t:0}$ also contains $r_{-1}$ and $r_{0}$.
$r_{-1}W_{K}^{1}w_{E}(v)$
$=\frac{\eta\alpha\alpha^{\prime}}{T}\sum_{t=1}^{T}\frac{P(t_{q}=t-1)}{t}$
×$\qquad\times\left(1-\frac{1+\mathbf{1}\{t=T\}}{T}\right)\left(1-\frac{1}{t}\right)$
$\quad-\frac{\eta\alpha\alpha^{\prime}}{T}\sum_{t=1}^{T}\frac{P(t_{q}=t)}{t^{2}}$
×$\qquad\times\left(\mathbf{1}\{t=T\}-\frac{2}{T}\right)$
$\quad-\frac{\eta\alpha\alpha^{\prime}}{VT^{2}}\sum_{k=1}^{V}\frac{P(t_{q}=T-1)}{t}\left(1-\frac{2}{T}\right),$
and
$r_{0}W_{K}^{1}w_{E}(v)$
$=\frac{\eta\alpha\alpha^{\prime}}{T}\sum_{t=1}^{T}\frac{P(t_{q}=t)}{t}$
×$\qquad\times\left(\mathbf{1}\{t=T\}-\frac{2}{T}\right)\left(1-\frac{1}{t}\right)$
$\quad-\frac{\eta\alpha\alpha^{\prime}}{T}\sum_{t=1}^{T}\frac{P(t_{q}=t-1)}{t^{2}}$
×$\qquad\times\left(1-\frac{1+\mathbf{1}\{t=T\}}{T}\right)$
$\quad+\frac{\eta\alpha\alpha^{\prime}}{VT^{2}}\sum_{k=1}^{V}P(t_{q}=T-1)\left(1-\frac{2}{T}\right).$
Also, for any $j\in\mathcal{V}$, we have
$w_{E}(j)W_{K}^{1}w_{E}(v)=O\left(\frac{1}{V}\right).$
The proof is finished. ∎
Appendix C Other theoretical proofs
Lemma 2 (Lemma 2, Bietti et al. [4]).
Given a data distribution $p$ over pairs $(x,y)\in\mathbb{R}^{d}\times[V]$ and a fixed matrix $W_{U}\in\mathbb{R}^{d\times d}$, minimizing the loss $L(W)=\mathbb{E}_{(x,y)\sim p}\left[\ell\left(y,W_{U}Wx\right)\right]$ gives the following gradient:
$\nabla_{W}L(W)=\sum_{v=1}^{V}p(y=v)w_{U}(v)(\hat{\mu}_{v}-\mu_{v})^{\top},$
where $\mu_{k}=\mathbb{E}[x\mid y=v]$ and $\hat{\mu}_{v}=\mathbb{E}_{x}[\frac{\hat{p}_{W}(v\mid x)}{p(y=v)}x]$, and $\hat{p}_{W}(v\mid x)$ is the probability of generating $v$.
Lemma 3.
Consider a two-layer attention only transformer that meets the initialization condition in Sec. B.2, and the input sequence of length $T$ ends with the trigger token $z_{T}=q$ at its second occurrence. Even after the gradient descent step of $W_{O}^{2}$ and $W_{K}^{2}$, the prediction $\hat{p}(v\mid z_{1:T})=1/V$ for any $v\in\mathcal{V}$.
Proof.
From the way the transformer is initialized, we have $W_{K}^{1}=O_{d\times d}$. Therefore, the output of the first attention block for any $t$ is:
$x_{t}^{(1)}=w_{E}(z_{t})+\frac{1}{t}\sum_{s=1}^{t}\Phi_{1}w_{E}(z_{s}).$
In the second layer, the key matrix is trained such that $(\Phi_{1}w_{E}(z_{t}))^{\top}W_{K}^{2}w_{E}(z_{T})=\alpha^{\prime}\mathbf{1}\{z_{t}=z_{T}\}$. Thus, we have
$\sigma\left((x_{1:T}^{(1)})^{\top}(W_{K}^{2})^{\top}W_{Q}^{2}x_{T}^{(1)}\right)$
$=\sigma\left(\begin{pmatrix}\left(w_{E}(z_{1})+\Phi_{1}w_{E}(z_{1})\right)^{\top}\\ \left(w_{E}(z_{2})+\sum_{i=1}^{2}\frac{\Phi_{1}w_{E}(z_{i})}{2}\right)^{\top}\\ \vdots\\ \left(w_{E}(z_{T-1})+\sum_{i=1}^{T-1}\frac{\Phi_{1}w_{E}(z_{i})}{T-1}\right)^{\top}\\ \left(w_{E}(z_{T})+\sum_{i=1}^{T}\frac{\Phi_{1}w_{E}(z_{i})}{T}\right)^{\top}\end{pmatrix}\right.$
×$\qquad\qquad\times\left.\left(\sum_{k\in Q}(\Phi_{1}w_{E}(k))w_{E}(k)^{\top}\right)Ix_{T}^{(1)}\right)$
$=\sigma\left(\begin{pmatrix}\left(w_{E}(z_{1})+\Phi_{1}w_{E}(z_{1})\right)^{\top}\\ \left(w_{E}(z_{2})+\sum_{i=1}^{2}\frac{\Phi_{1}w_{E}(z_{i})}{2}\right)^{\top}\\ \vdots\\ \left(w_{E}(z_{t-1})+\sum_{i=1}^{t-1}\frac{\Phi_{1}w_{E}(z_{i})}{t-1}\right)^{\top}\\ \left(w_{E}(z_{t})+\sum_{i=1}^{t}\frac{\Phi_{1}w_{E}(z_{i})}{t}\right)^{\top}\end{pmatrix}\right.$
×$\qquad\qquad\left.\times\Phi_{1}w_{E}(q)\right)$
$=\sigma\left(\frac{\eta\tau}{TV}\begin{pmatrix}0\\ \vdots\\ 0\\ 1/(t_{o}-1)\\ 1/t_{o}\\ \vdots\\ 1/(T-1)\\ 1/T\end{pmatrix}\right).$
When $V$ is large enough, the attention is spread across the whole sequence evenly, and we obtain
$x_{T}^{(2)}$ (32)
$\approx\frac{1}{T}\sum_{t=1}^{T}W_{O}^{2}W_{V}^{2}x_{t}^{(1)}+x_{T}^{(1)}$ (33)
$=\frac{1}{T}\sum_{t=1}^{T}W_{O}^{2}W_{V}^{2}\left(w_{E}(z_{t})+\frac{1}{t}\sum_{s=1}^{t}\Phi_{1}w_{E}(z_{s})\right)$
$\quad+w_{E}(z_{T})+\frac{1}{T}\sum_{s=1}^{T}\Phi_{1}w_{E}(z_{s})$
$=_{w_{U}}\frac{\alpha}{T}\sum_{t=1}^{T}w_{U}(z_{t}),$
where we focus on the terms represented by $w_{U}(v)$ for vocabulary $v$ by $=_{w_{U}}$. Taking the expectation over $z_{1:T}$ while keeping the trigger token $z_{T}=z_{t_{o}-1}=q$ fixed yields
$\mathbb{E}[x_{T}^{(2)}]$
$=_{w_{U}}\sum_{j=1}^{V}\mathbb{E}\left[x_{T}^{(2)}\mid q=j\right]$
$=\frac{\alpha}{T}\sum_{j=1}^{V}\left(\frac{T-2}{V-1}\sum_{\begin{subarray}{c}1\leq v\leq V\\ v\neq j\end{subarray}}w_{U}(v)+2w_{U}(j)\right)$
The coefficients of $w_{U}(v)$ are the same for any $v\in\mathcal{V}$ and this concludes the proof. ∎
Lemma 4.
Consider a two-layer attention only transformer that meets the initialization condition in Sec. B.2, and the input sequence of length $T$ ends with the trigger token $z_{T}=q$ at its second occurrence. After the gradient descent step of $W_{O}^{2}$, the prediction $\hat{p}(v\mid z_{1:T})=1/V$ for any $v\in\mathcal{V}$.
Proof.
Since $W_{K}^{2}=O_{d\times d}$, it is guaranteed that we have Eq. 33 as an equation. Thus, the result follows directly from Lemma  3. ∎
Lemma 5.
Suppose the loss function is given by
$L(W)$ $=\mathbb{E}_{(X,y)}[l(y,\xi(X))]$
$\xi(X)$ $=W_{U}\Phi_{2}X\bar{\sigma}(Z(W)^{\top}W_{2}x_{T}),$
where $Z(W)=(z_{1}(W),\dots,z_{T}(W))$ with $z_{t}(W)=\sum_{s=1}^{t}\Phi_{1}x_{s}\sigma(\left(x_{1:t}+R_{1-t:0}\right)^{\top}Wx_{t})_{s}$, and $\bar{\sigma}(u_{1:T})_{t}=\frac{1}{T}(1+u_{t}-\frac{1}{T}\sum_{s=1}^{T}u_{s})$ is the linearization of the softmax function around 00. Then the gradient at $W=0$ has the following form.
$\nabla_{W}L(W)$
$=\sum_{i=1}^{N}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}w_{U}(i)^{\top}\Phi_{2}x_{t}\right.$
×$\quad\left.\times\frac{1}{t}\sum_{s=1}^{t}(\Phi_{1}x_{s})^{\top}W_{K}^{2}x_{T}q_{s,t}x_{t}^{\top}\mid y=i\right]$
$\quad-\sum_{i=1}^{N}\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}w_{U}(i)^{\top}\Phi_{2}x_{t}\right.$
×$\quad\left.\times\frac{1}{t}\sum_{s=1}^{t}(\Phi_{1}x_{s})^{\top}W_{K}^{2}x_{T}q_{s,t}x_{t}^{\top}\right]$
$\quad-\sum_{i=1}^{N}\mathbb{E}\left[w_{U}(i)^{\top}\Phi_{2}\bar{x}_{1:T}\right.$
×$\quad\left.\times\frac{1}{T}\sum_{t=1}^{T}\frac{1}{t}\sum_{s=1}^{t}(\Phi_{1}x_{s})^{\top}W_{K}^{2}x_{T}q_{s,t}x_{t}^{\top}\mid y=i\right]$
$\quad+\sum_{i=1}^{N}\mathbb{E}\left[w_{U}(i)^{\top}\Phi_{2}\bar{x}_{1:T}\right.$
×$\quad\left.\times\frac{1}{T}\sum_{t=1}^{T}\frac{1}{t}\sum_{s=1}^{t}(\Phi_{1}x_{s})^{\top}W_{K}^{2}x_{T}q_{s,t}x_{t}^{\top}\right],$
where $q_{s,t}=(x_{s}+r_{s-t})-(\bar{x}_{1:t}-\bar{r}_{1-t:0})$ and $\bar{x}_{1:t}=\frac{1}{t}\sum_{s=1}^{t}x_{s}$.
Proof.
The result is obtained by replacing $p_{1:t}^{\top}$ with $\left(x_{1:t}+R_{1-t:0}\right)^{\top}$ and $p_{t}$ with $x_{t}$ in $z_{t}$ from the proof of Lemma 5 from Bietti et al. [4]. ∎
It is worth noting that $x_{i}=x^{(0)}_{i}=w_{E}(z_{i})$ and that if we directly follows the transformer architecture the vector in the linearized version of softmax $\bar{\sigma}$ should be $Z^{\prime}(W)^{\top}W_{2}z^{\prime}_{T}(W)$, where $Z^{\prime}(W)=(z^{\prime}_{1}(W),\dots,z^{\prime}_{T}(W))$ with $z^{\prime}_{t}(W)=x_{t}+\sum_{s=1}^{t}\Phi_{1}x_{s}\sigma(\left(x_{1:t}+R_{1-t:0}\right)^{\top}Wx_{t})_{s}$. Since $W_{K}^{2}$ is learned so that it only cares the ordered pair of $\Phi_{1}w_{E}(z_{t})$ and $w_{E}(z_{t})$, the loss function is simplified and ignores the other terms. This also applies to $\xi(X)$, where $\xi^{\prime}(X)=W_{U}\Phi_{2}Z^{\prime}\bar{\sigma}(Z^{\prime}(W)^{\top}W_{2}z^{\prime}_{T})$ is the authentic form of the output. Since we know $W_{O}^{2}$ (and therefore $\Phi_{2}$) is already trained, we can simplify the expression.
C.1 Supporting lemmas
This section provides useful lemmas to show the learning of matrices, especially for $W_{K}^{1}$.
Lemma 6.
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. For any $t\in[1,T]$, and any vocabulary $k\in\mathcal{V}$, we have
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=t-1]=1.$
Proof.
Since the trigger token is $z_{t_{q}}=z_{t-1}$, the next token $z_{t}$ is the output token. The condition $y=k$ gives us $z_{t}=k$. Therefore, $\mathbf{1}\{z_{t}=k\}=1$ and the equation holds. ∎
Lemma 7.
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. Fix some $t\in[1,T]\subset\mathbb{N}$. For any $s\in[1,t-2]\cup\{t\}$, and vocabulary $k\in\mathcal{V}$, we have
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s]$
=$=\begin{cases}1&\text{if $s=t$ and $t=T-1$},\\ \frac{1}{V-1}&\text{if $s\leq t-2$ and $t\neq T$},\\ 0&\text{otherwise},\end{cases}$ (34)
Proof.
First, consider the case $s=t$. Under the condition $t_{q}=s$, we have $t_{q}=s=t$, and thus, $z_{t}$ is the trigger token. The only possible case allowing the same token for trigger and output is when $t_{q}=T-1$. This yields $z_{T-1}=z_{T}=k$ because $z_{T}$ not only works as trigger, but also output. Therefore, when $s=t$, we have
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s]$
=$=\begin{cases}1&\text{if $t=T-1$},\\ 0&\text{otherwise}.\end{cases}$ (35)
Next, we focus on $s\leq t-2$. Suppose we are given $t=T$. The token $z_{t}$ is then the second trigger token, and $\mathbf{1}\{z_{t}=k\}=1$ implies that $z_{t_{q}}=k$. However, since the first trigger position is $t_{q}=s\leq t-2$ and the output token is also $y=k$, $z_{T}(=k)$ cannot be the second trigger occurrence. Hence, it suffices to consider $t\neq T$. We now have $s\leq t-2<T-2$. Assuming $z_{t_{q}}=k$ will also lead to contradiction because the trigger token $k$ appears more than twice at positions $s,s+1$ and $T$. With this in mind, we obtain
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s]$
$=P(z_{t}=k\mid y=k,t_{q}=s)$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}P(z_{t}=k,z_{t_{q}}=q\mid y=k,t_{q}=s)$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}P(z_{t}=k\mid y=k,t_{q}=s,z_{t_{q}}=q)$
×$\quad\times P(z_{t_{q}}=q\mid y=k,t_{q}=s)$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{1}{V-1}\frac{1}{V-1}$
$=\frac{1}{V-1}.$ (36)
Here we use $P(z_{t_{q}}=q\mid y=k,t_{q}=s)=1/(V-1)$ because the trigger token is uniformly sampled from $\mathcal{V}\setminus\{k\}$. Also, we have $P(z_{t}=k\mid y=k,t_{q}=s,z_{t_{q}}=q)=1/(V-1)$ because $z_{t}$ for $t\neq T$ is neither the trigger token nor the output token, and thus, sampled from $\mathcal{V}\setminus\{q\}$. These give us the following when $s\leq t-2$:
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s]=\begin{cases}\frac{1}{V-1}&\text{if $t\neq T$},\\ 0&\text{otherwise}.\end{cases}$ (37)
Combining Eq. C.1 and 37, we obtain the result. ∎
Lemma 8.
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. For any vocabulary $k\in\mathcal{V}$, we have
$\mathbb{E}\left[\frac{1}{T}\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\right.$ (38)
×$\quad\left.\times\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=t-1\right]$
=$=\begin{cases}\frac{2}{T}&\text{if $t=T$},\\ \frac{1}{T}\left(1+\frac{T-3}{V-1}\right)&\text{otherwise}.\end{cases}$ (39)
Proof.
By the linearity of expectation, we can decompose the expectation as follows:
$\mathbb{E}\left[\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$=\mathbb{E}[\mathbf{1}\{z_{t_{q}}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]$
$\quad+\mathbb{E}[\mathbf{1}\{z_{t_{q}+1}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]$
$\quad+\mathbf{1}\{t\neq T\}\mathbb{E}[\mathbf{1}\{z_{T}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]$
$\quad+\mathbb{E}\left[\left(\sum_{t^{\prime}\neq t_{q},t_{q}+1,T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\mathbf{1}\{z_{t}=k\}\mid\cdot\right],$
where we abbreviate the conditions of the expectations to enhance readability.
Note that we always have $\mathbf{1}\{z_{t}=k\}=1$ when $t_{q}=t-1$ and $y=k$. In the case of $t=T$, the trigger token appears at positions $T-1$ and $T$ given $t_{q}=t-1$. With the condition $y=k$, it follows that
$\mathbf{1}\{z_{t_{q}}=k\}\mathbf{1}\{z_{t}=k\}=1,$
$\mathbf{1}\{z_{t_{q}+1}=k\}\mathbf{1}\{z_{t}=k\}=1.$
In the above, we point out that we have $t_{q}+1=T$ in this situation. Additionally, we get
$\mathbb{E}\left[\left(\sum_{t^{\prime}\neq t_{q},t_{q}+1,T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\mathbf{1}\{z_{t}=k\}\mid\cdot\right]=0,$
where the final equation follows from the fact that $z_{t^{\prime}}$ is not a trigger token $k$.
For the case where $t\neq T$, the trigger token is not $k$ because otherwise there will be more than two triggers in the input sequence at positions $t_{q},t_{q}+1$ and $T$. Thus, we have
$\mathbb{E}[\mathbf{1}\{z_{t_{q}}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]=0,$
$\mathbb{E}[\mathbf{1}\{z_{t_{q}+1}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]=1,$
$\mathbb{E}[\mathbf{1}\{z_{T}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]=0.$
Following the same argument in Eq. 36, we also have
$\mathbb{E}\left[\left(\sum_{t^{\prime}\neq t_{q},t_{q}+1,T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$=\mathbb{E}\left[\sum_{t^{\prime}\neq t_{q},t_{q}+1,T}\mathbf{1}\{z_{t^{\prime}}=k\}\mid\cdot\right]$
$=\frac{T-3}{V-1}.$
We prove the statement by combining these results. ∎
Lemma 9.
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. Fix some $t\in[1,T]\subset\mathbb{N}$. For any $s\in[1,t-2]\cup\{t\}$, and vocabulary $k\in\mathcal{V}$,
$\mathbb{E}\left[\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s\right]$
=$=\begin{cases}2&\text{if $s=t$ and $t=T-1$},\\ \frac{2}{V-1}+\frac{T-4}{(V-1)^{2}}&\text{if $s\leq t-2$ and $t\neq T$},\\ 0&\text{otherwise}.\end{cases}$ (40)
Proof.
First, we look at the scenario $s=t$. Since $t_{q}=s=t$, the input sequence $z_{1:T}$ will contain more than two trigger tokens unless $t=T-1$, which does not satisfy the condition of input sequence, we examine the case of $t=T-1$. This condition ensures that $\mathbf{1}\{z_{T-1}=k\}=\mathbf{1}\{z_{T}=k\}=1$, and therefore, we can use the instance $t=T$ from Lemma 8.
$\mathbb{E}\left[\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\mathbf{1}\{z_{t}=k\}\mid\cdot\right]=2.$
Next, we require $s\leq t-2$. It is impossible to have $t=T$ because this violates the input sequence condition, where the trigger token only appears twice. Now, by the linearity of expectation, we can decompose the expectation as follows:
$\mathbb{E}\left[\left(\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$=\mathbb{E}[\mathbf{1}\{z_{t_{q}}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]$
$\quad+\mathbb{E}[\mathbf{1}\{z_{t_{q}+1}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]$
$\quad+\mathbb{E}[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]$
$\quad+\mathbb{E}[\mathbf{1}\{z_{T}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]$
$\quad+\mathbb{E}\left[\left(\sum_{t^{\prime}\neq t_{q},t_{q}+1,t,T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\mathbf{1}\{z_{t}=k\}\mid\cdot\right].$
Since $t_{q}=s\leq t-2$ and $t\neq T$, the token $z_{t}$ is neither trigger nor output. This yields the following:
$\mathbb{E}[\mathbf{1}\{z_{t_{q}}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]=0,$
$\mathbb{E}[\mathbf{1}\{z_{T}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]=0.$
Given $y$ equals $k$, we have $\mathbf{1}\{z_{t_{q}+1}=k\}=1$. Therefore, from Lemma 7, we get
$\mathbb{E}[\mathbf{1}\{z_{t_{q}+1}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]$
$=\mathbb{E}[\mathbf{1}\{z_{t}=k\}\mid\cdot]=\frac{1}{V-1},$
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{z_{t}=k\}\mid\cdot]$
$=\mathbb{E}[\mathbf{1}\{z_{t}=k\}\mid\cdot]=\frac{1}{V-1}.$
Finally, for $t^{\prime}\neq t_{q},t_{q}+1,t$ and $T$, the tokens $z_{t^{\prime}}$ and $z_{t}$ are independent. With Lemma 7, we obtain
$\mathbb{E}\left[\left(\sum_{t^{\prime}\neq t_{q},t_{q}+1,t,T}\mathbf{1}\{z_{t^{\prime}}=k\}\right)\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$=\frac{T-4}{(V-1)^{2}}.$
Integrating the equations above allows us to arrive at the conclusion. ∎
Lemma 10.
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. For any vocabulary $k\in\mathcal{V}$, and $t\in[1,T]\subset\mathbb{N}$, we have
$\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}x_{t_{q}}\mid y=k\right]$
$=P(t_{q}=t-1)\mathbf{1}\{t=T\}w_{E}(k)$
$\quad+P(t_{q}=t)\mathbf{1}\{t=T-1\}w_{E}(k)+O\left(\frac{1}{V}\right).$
Proof.
We split $\mathbf{1}\{t_{q}\leq t\}$ into $\mathbf{1}\{t_{q}=t-1\}+\mathbf{1}\{t_{q}=t\}+\mathbf{1}\{t_{q}\leq t-2\}$. Then, the expectation can be decomposed as:
$\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}x_{t_{q}}\mid y=k\right]$
$=\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}=t-1\}x_{t_{q}}\mid y=k\right]$
$\quad+\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}=t\}x_{t_{q}}\mid y=k\right]$
$\quad+\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t-2\}x_{t_{q}}\mid y=k\right].$
Note that when $t_{q}=t-1$, $z_{t}$ is the output token. In addition, if we have $t=T$, $z_{t}$ works as the second output token as well, which means that $z_{t}=k$ and $x_{t_{q}}=w_{E}(k)$. If we find $t\neq T$, the trigger token is not $k$. Thus, we derive
$\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}=t-1\}x_{t_{q}}\mid y=k\right]$
$=P(t_{q}=t-1)$
×$\qquad\times\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}x_{t_{q}}\mid y=k,t_{q}=t-1\right]$
$=P(t_{q}=t-1)\mathbf{1}\{t=T\}w_{E}(k)$
$\quad+P(t_{q}=t-1)\mathbf{1}\{t\neq T\}\sum_{j\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(j)}{V-1}.$
Next, we discuss $t_{q}=t$. Since $t_{q}$ is the first trigger position, we always observe that $t_{q}\leq T-1$. If two trigger tokens are adjacent to each other, i.e., $t_{q}=T-1$, we can state that $z_{T-1}=z_{T}=k$ because of the equation $y=k$. Otherwise, the trigger token must belong to $\mathcal{V}\setminus\{k\}$, and $\mathbf{1}\{z_{t}=k\}=0$. This gives
$\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}=t\}x_{t_{q}}\mid y=k\right]$
$=P(t_{q}=t-1)\mathbf{1}\{t=T\}w_{E}(k).$
For $t_{q}\leq t-2$, we again analyze the cases of $t=T$ and $t\neq T$. Provided that $t=T$, $\mathbf{1}\{z_{t}=k\}=1$ implies that the trigger token is decided to be $k$, while the output token is also $k$. This leads to more than two occurrences of trigger token, which violates the input sequence condition. Now, we assume $t\leq T-1$. This time, the trigger token $z_{t_{q}}$ is not $k$, and $z_{t}$ is not a trigger token, which yields
$\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t-2\}x_{t_{q}}\mid y=k\right]$
$=P(t_{q}\leq t-2)\mathbb{E}[w_{E}(z_{t_{q}})\cdot\mathbb{E}[\mathbf{1}\{z_{t}=k\}$
$\quad\mid z_{t_{q}},y=k,t_{q}\leq t-2]\mid y=k,t_{q}\leq t-2]$
$=P(t_{q}\leq t-2)\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(q)}{V-1}$
×$\quad\times\mathbb{E}[\mathbf{1}\{z_{t}=k\}\mid z_{t_{q}}=q,y=k,t_{q}\leq t-2]$
$=P(t_{q}\leq t-2)\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(q)}{(V-1)^{2}}.$
This concludes the proof. ∎
Lemma 11.
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. For any vocabulary $k\in\mathcal{V}$, and $t\in[1,T]\subset\mathbb{N}$, we have
$\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}\leq t\}\bar{x}_{1:t}\mid y=k\right]$
$=\frac{P(t_{q}=t-1)}{t}(1+\mathbf{1}\{t=T\})w_{E}(k)$
$\quad+\frac{\mathbf{1}\{t=T-1\}P(t_{q}=t)}{t}w_{E}(k)+O\left(\frac{1}{V}\right).$ (41)
Proof.
As in the proof of Lemma 10, we analyze each scenario depending on the value of $t_{q}$. Firstly, let us consider $t_{q}=t-1$. This leads to the fact that $z_{t}$ is the output token. Given $y=k$, we have $z_{t}=k$.
$\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}\mathbf{1}\{t_{q}=t-1\}\bar{x}_{1:t}\mid y=k\right]$
$=P(t_{q}=t-1)\mathbb{E}\left[\bar{x}_{1:t}\mid y=k,t_{q}=t-1\right].$
In the event that $t=T$ is valid, $x_{t_{q}}=x_{T}=w_{E}(k)$ holds and the other $T-2$ tokens are not $k$. Thus, it gives
$\mathbb{E}\left[\bar{x}_{1:T}\mid y=k\right]$
$=\frac{1}{T}\left\{2w_{E}(k)+(T-2)\sum_{j\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(j)}{V-1}\right\}.$
When $t\neq T$ holds, we can say that $x_{t_{q}}\neq w_{E}(k)$, while $x_{t_{q}+1}=w_{E}(k)$ because of $y=k$. This gives us
$\mathbb{E}[x_{t_{q}}\mid y=k,t_{q}=t-1]=\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(j)}{V-1},$
$\mathbb{E}[x_{t_{q}+1}\mid y=k,t_{q}=t-1]=w_{E}(k).$
Also, we can derive the following equations for $i\in[1,t]\setminus\{t_{q},t_{q}+1\}$:
$\mathbb{E}[x_{i}\mid y=k,t_{q}=t-1]$
$=\sum_{j\in\mathcal{V}}P(z_{i}=j\mid\cdot)w_{E}(j)$
$=\sum_{q=1}^{V}\sum_{j=1}^{V}P(z_{i}=j,z_{t_{q}}=q\mid\cdot)w_{E}(j)$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}\sum_{j\in\mathcal{V}\setminus\{q\}}P(z_{i}=j,z_{t_{q}}=q\mid\cdot)w_{E}(j)$
$=\sum_{\begin{subarray}{c}q\in\mathcal{V}\setminus\{k\}\\ j\in\mathcal{V}\setminus\{q\}\end{subarray}}P(z_{i}=j,\mid z_{t_{q}}=q,\cdot)$
×$\qquad\times P(z_{t_{q}}=q\mid\cdot)w_{E}(j)$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}\sum_{j\in\mathcal{V}\setminus\{q\}}\frac{w_{E}(j)}{(V-1)^{2}}.$
In the above, we use the fact that $P(x_{i}=j,x_{t_{q}}=q)=0$ if $j=q$ or $q=k$ holds. To summarize, we have
$\mathbb{E}\left[\bar{x}_{1:t}\mid y=k,t_{q}=t-1\right]$
=$=\begin{cases}\frac{2w_{E}(k)}{t}+\frac{t-2}{t}\sum_{j\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(j)}{V-1}&\text{if $t=T$},\\ \frac{1}{t}\left\{w_{E}(k)+\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(q)}{V-1}\right\}\\ \quad+\frac{t-2}{t}\left\{\sum_{\begin{subarray}{c}q\in\mathcal{V}\setminus\{k\}\\ j\in\mathcal{V}\setminus\{q\}\end{subarray}}\frac{w_{E}(j)}{(V-1)^{2}}\right\}&\text{otherwise}.\end{cases}$ (42)
=$=\begin{cases}\frac{2}{T}w_{E}(k)+O\left(\frac{1}{V}\right)&\text{if $t=T$},\\ \frac{1}{t}w_{E}(k)+O\left(\frac{1}{V}\right)&\text{otherwise}.\end{cases}$
Next, assume $t_{q}=t$. In this case, $t=T$ implies $t_{q}=T$ which is a contradiction. Furthermore, $t_{q}=t<T-1$ implies $t_{q}<t_{q+1}<T$. Under this condition, we cannot have $z_{t}=k$, because it produces three appearances of the trigger token. Hence, it suffices to consider $t=T-1$. If $t_{q}=t=T-1$ is true, we can state that $x_{t}=k$, leading to the formulation below:
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}\bar{x}_{1:t}\mid y=k,t_{q}=t]$
$=\mathbb{E}[\bar{x}_{1:t}\mid y=k,t_{q}=t]$
$=\frac{1}{t}\left\{1w_{E}(k)+(T-2)\sum_{\begin{subarray}{c}1\leq j\leq V\\ j\neq k\end{subarray}}\frac{w_{E}(j)}{V-1}\right\}$
$=\frac{1}{T-1}w_{E}(k)+O\left(\frac{1}{V}\right),$
where the second equation follows from the same argument in the instance $t=T$ in Eq. 42, but this time $x_{T}$ is not contained in $x_{1:t}$.
Lastly, we focus on $t_{q}=s\leq t-2$. We argue that the trigger token $z_{t_{q}}=z_{T}$ is not $k$, due to the condition of the number of trigger token in the input sequence. This means that $t\neq T$ and $z_{t}$ does not serve as a trigger token. Now, we split $\bar{x}_{1:t}$ into $\frac{1}{t}(x_{t_{q}}+x_{t_{q}+1}+x_{t}+\sum_{t^{\prime}}x_{t^{\prime}})$ and obtain
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}x_{t_{q}}\mid y=k,t_{q}=s]$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}P(z_{t}=k,z_{t_{q}}=q\mid\cdot)w_{E}(q)$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(q)}{(V-1)^{2}},$
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}x_{t_{q}+1}\mid y=k,t_{q}=s]$
$=P(z_{t}=k\mid\cdot)w_{E}(k)$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}P(z_{t}=k,z_{t_{q}}=q\mid\cdot)w_{E}(k)$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(k)}{(V-1)^{2}},$
$=\frac{w_{E}(k)}{V-1},$
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}x_{t}\mid y=k,t_{q}=s]$
$=P(z_{t}=k\mid y=k,t_{q}=s)w_{E}(k)$
$=\frac{w_{E}(k)}{V-1},$
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}x_{t^{\prime}}\mid y=k,t_{q}=s]$
$=\sum_{\begin{subarray}{c}q\in\mathcal{V}\setminus\{k\}\\ j\in\mathcal{V}\setminus\{q\}\end{subarray}}P(z_{t}=k,z_{t^{\prime}}=j,z_{t_{q}}=q\mid\cdot)w_{E}(j)$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}P(z_{t_{q}}=q\mid\cdot)$
×$\quad\times\sum_{j\in\mathcal{V}\setminus\{q\}}P(z_{t}=k,z_{t^{\prime}}=j\mid z_{t_{q}}=q,\cdot)w_{E}(j)$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{1}{V-1}\sum_{j\in\mathcal{V}\setminus\{q\}}\frac{w_{E}(j)}{(V-1)^{2}}.$
In short, we have
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}\bar{x}_{1:t}\mid y=k,t_{q}=s]=O\left(\frac{1}{V}\right).$
Summing up all the related terms leads to Eq. 41, which finishes the proof. ∎
Lemma 12.
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. For any vocabulary $k\in\mathcal{V}$, and $t\in[1,T]\subset\mathbb{N}$, we have
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right.$
×$\quad\left.\times\mathbf{1}\{t_{q}\leq t\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k\right]$
$=\frac{2w_{E}(k)}{T}\mathbf{1}\{t=T\}P(t_{q}=t-1)$
$\quad+\frac{2w_{E}(k)}{T}\mathbf{1}\{t=T-1\}P(t_{q}=t)$
$\quad+\frac{1}{T}\cdot O\left(\frac{1}{V}\right).$ (43)
Proof.
We consider the following cases:
Case 1: $t_{q}=t-1$.
If $t=T$ holds, this gives $z_{t_{q}}=z_{T}=y=k$ and $\mathbf{1}\{z_{t-1}=k\}=\mathbf{1}\{z_{t}=k\}=1$. Thus, we have
$\mathbb{E}\left[\mathbf{1}\{z_{t-1}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=t-1\right]$
$=\mathbb{E}\left[x_{t_{q}}\mid y=k,t_{q}=t-1\right]=w_{E}(k),$
$\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=t-1\right]$
$=\mathbb{E}\left[x_{t_{q}}\mid y=k,t_{q}=t-1\right]=w_{E}(k).$
Since $z_{t^{\prime}}$ for $t^{\prime}\neq t_{q},t_{q}+1$ is not a trigger token, it follows that $\mathbf{1}\{z_{t^{\prime}}=k\}=0$, and we have
$\mathbb{E}\left[\sum_{t^{\prime}\neq t-1,t}\mathbf{1}\{z_{t^{\prime}}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]=0.$
Otherwise, when $t\neq T$, $z_{t_{q}}=z_{T}$ is not $k$, yielding
$\mathbb{E}\left[\mathbf{1}\{z_{t-1}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]=0.$
$\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]=0.$
Also, $z_{t_{q}}$ is sampled from $\mathcal{V}\setminus\{k\}$, we get
$\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=t-1\right]$
$=$ $\mathbb{E}\left[x_{t_{q}}\mid y=k,t_{q}=t-1\right]$
$=$ $\sum_{j\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(j)}{V-1}.$
For the rest of $t^{\prime}$, we note that $z_{t_{q}}\neq k$ and obtain
$\mathbb{E}\left[\sum_{t^{\prime}\neq t,t-1,T}\mathbf{1}\{z_{t^{\prime}}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$=\sum_{\begin{subarray}{c}t^{\prime}\\ q\in\mathcal{V}\setminus\{k\}\end{subarray}}P(z_{t^{\prime}}=k,z_{t_{q}}=q\mid\cdot)w_{E}(q)$
$=(T-3)\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(q)}{(V-1)^{2}}.$ (44)
In summary, we find that
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right.$
×$\quad\left.\times\mathbf{1}\{t_{q}=t-1\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k\right]$
$=$ $\frac{P(t_{q}=t-1)}{T}\mathbf{1}\{t=T\}2w_{E}(k)$
$\quad+\frac{P(t_{q}=t-1)}{T}\mathbf{1}\{t\neq T\}O\left(\frac{1}{V}\right).$ (45)
Case 2: $t_{q}=t$.
For this scenario, $\mathbf{1}\{z_{t}=k\}=1$ holds only when $t_{q}=t=T-1$. Thus, we can state that
$\mathbb{E}[\mathbf{1}\{z_{t_{q}}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=t]$
$=\mathbb{E}[x_{t_{q}}\mid y=k,t_{q}=t]=w_{E}(k),$
$\mathbb{E}[\mathbf{1}\{z_{T}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=t]$
$=\mathbb{E}[x_{t_{q}}\mid y=k,t_{q}=t]=w_{E}(k).$
For the other $t^{\prime}$, we have $\mathbf{1}\{z_{t^{\prime}}=k\}=0$ because these tokens are not a trigger token. This means that
$\mathbb{E}\left[\sum_{t^{\prime}\neq t,T}\mathbf{1}\{z_{t^{\prime}}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=t\right]$
$=0.$
In short, we have
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\mathbf{1}\{t_{q}=t\}\right.$
×$\qquad\left.\times x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k\right]$
$=\frac{\mathbf{1}\{t=T-1\}P(t_{q}=t)}{T}2w_{E}(k).$ (46)
Case 3: $t_{q}\leq t-2$.
We first let $t=T$. If $\mathbf{1}\{z_{t}=k\}=1$ holds, then the trigger token becomes $k$, but we find that $z_{t_{q}}=z_{t_{q}+1}=z_{T}=k$, meaning the trigger token appears more than two times. Therefore, we only consider $t\neq T$. Given $z_{t_{q}}(=z_{T})=k$, it also fails to satisfy the required conditions for the sequence, hence it follows that
$\mathbb{E}[\mathbf{1}\{z_{s}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s]=0,$
$\mathbb{E}[\mathbf{1}\{z_{T}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s]=0,$
where $21\leq s\leq t-$ 2.
Similarly to Eq. 44, we can derive that
$\mathbb{E}[\mathbf{1}\{z_{t}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s]$
$=\mathbb{E}[x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s]$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(q)}{(V-1)^{2}},$
$\mathbb{E}[\mathbf{1}\{z_{s+1}=k\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s]$
$=\mathbb{E}[x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s]$
$=\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(q)}{(V-1)^{2}}.$
In addition, we have
$\mathbb{E}\left[\sum_{t^{\prime}\neq s,s+1,t,T}\mathbf{1}\{z_{t^{\prime}}=k\}x_{t_{q}}\right.$
×$\quad\left.\times\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s\right]$
$=\sum_{\begin{subarray}{c}t^{\prime}\\ q\in\mathcal{V}\setminus\{k\}\end{subarray}}P(z_{t^{\prime}}=k,z_{t}=k,z_{t_{q}}=q\mid\cdot)w_{E}(q)$
$=\sum_{\begin{subarray}{c}t^{\prime}\\ q\in\mathcal{V}\setminus\{k\}\end{subarray}}P(z_{t^{\prime}}=k,z_{t}=k,\mid z_{t_{q}}=q,\cdot)$
×$\quad\times P(z_{t_{q}}=q\mid y=k,t_{q}=s)w_{E}(q)$
$=(T-4)\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(q)}{(V-1)^{3}}.$
These equations give us that for $21\leq s\leq t-$ 2,
$\mathbb{E}\left[\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\mathbf{1}\{t_{q}=s\}x_{t_{q}}\right.$ (47)
×$\quad\left.\times\mathbf{1}\{z_{t}=k\}\mid y=k\right]$
$=O\left(\frac{1}{V^{2}}\right).$ (48)
Gathering all the equations of $1\leq t_{q}\leq t$ leads to Eq. 43 and we have proven the statement. ∎
Lemma 13.
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. For any vocabulary $k\in\mathcal{V}$, and $t\in[1,T]\subset\mathbb{N}$, we have
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right.$
×$\quad\left.\times\mathbf{1}\{t_{q}\leq t\}\bar{x}_{1:t}\mathbf{1}\{z_{t}=k\}\mid y=k\right]$
$=\frac{(4\cdot\mathbf{1}\{t=T\}+\mathbf{1}\{t\neq T\})}{Tt}$
×$\quad\times P(t_{q}=t-1)\cdot w_{E}(k)$
$\quad+\frac{2\cdot\mathbf{1}\{t=T-1\}P(t_{q}=t)}{Tt}w_{E}(k)$
$\quad+O\left(\frac{1}{VT}\right).$ (49)
Proof.
Considering the linearity of expectation, we focus on each term $x_{i}$ contained in $\bar{x}_{1:t}$. We consider the following cases:
Case 1: $t_{q}=t-1$.
We already have the result as to $x_{t_{q}}$ in Eq. C.1:
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right.$
×$\quad\left.\times\mathbf{1}\{t_{q}=t-1\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k\right]$
$=\frac{P(t_{q}=t-1)}{T}\mathbf{1}\{t=T\}2w_{E}(k)$
$\quad+\frac{P(t_{q}=t-1)}{T}\mathbf{1}\{t\neq T\}O\left(\frac{1}{V}\right).$
We can also use Lemma 8 for $x_{t}=k$, i.e.,
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}x_{t}\mathbf{1}\{z_{t}=k\}\mid y=k\right]$
$=\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\mid y=k,t_{q}=t-1\right]$
×$\quad\times P(t_{q}=t-1)w_{E}(k)$
=$=\begin{cases}\frac{P(t_{q}=t-1)}{T}2w_{E}(k)&\text{if $t=T$},\\ \frac{P(t_{q}=t-1)}{T}\left(1+\frac{T-3}{V-1}\right)w_{E}(k)&\text{otherwise}.\end{cases}$
Now we deal with $x_{i}$ for $21\leq i\leq t-$ 2. Note that $z_{i}\neq k$ for these $i$ under the conditions $t=T$ and $y=k$.
$\mathbb{E}\left[\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=t-1\right]$
$=\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$\quad+\mathbb{E}\left[\mathbf{1}\{z_{t_{q}}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$\quad+\mathbb{E}\left[\mathbf{1}\{z_{i}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$\quad+\sum_{t^{\prime}}\mathbb{E}\left[\mathbf{1}\{z_{i}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
=$=\begin{cases}\sum_{j\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(j)}{V-1}\\ \quad+\sum_{j\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(j)}{V-1}&\text{if $t=T$},\\ \sum_{j\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(j)}{V-1}+\frac{w_{E}(k)}{V-1}\\ +(T-3)\sum_{\begin{subarray}{c}q\in\mathcal{V}\setminus\{k\}\\ j\in\mathcal{V}\setminus\{q\}\end{subarray}}\frac{w_{E}(j)}{(V-1)^{3}}&\text{otherwise}\end{cases}$
$=O\left(\frac{1}{V}\right)$
Case 2: $t_{q}=t$.
For this scenario, we have seen that $t_{q}=t=T-1$ is the only possible way for the expectation to take non-zero value. For $x_{t_{q}}$, we can observe Eq. C.1 and obtain
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right.$
×$\quad\left.\times\mathbf{1}\{t_{q}=t\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k\right]$
$=\frac{\mathbf{1}\{t=T-1\}P(t_{q}=t)}{T}2w_{E}(k).$
The other terms related to $x_{i}$ for $11\leq i\leq t-$ 1 can be calculated
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=t\right]$
$=\frac{1}{T}\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$\quad+\frac{1}{T}\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$\quad+\frac{1}{T}\sum_{t^{\prime}\neq t}\mathbb{E}\left[\mathbf{1}\{z_{t^{\prime}}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid\cdot\right].$
$=\frac{1}{T}\left(\sum_{j\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(j)}{V-1}+\sum_{j\in\mathcal{V}\setminus\{k\}}\frac{w_{E}(j)}{V-1}+0\right)$
$=\frac{1}{T}\cdot O\left(\frac{1}{V}\right).$
We note that $\mathbf{1}\{z_{t}=k\}=1$ and $\mathbf{1}\{z_{t^{\prime}}=k\}=0$ for $t^{\prime}<t_{q}=t$ because of the input sequence condition.
Case 3: $t_{q}=s\leq t-2$.
Again, we have the result for $x_{t_{q}}$ in Eq. 47.
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right.$
×$\quad\left.\times\mathbf{1}\{t_{q}=s\}x_{t_{q}}\mathbf{1}\{z_{t}=k\}\mid y=k\right]$
$=O\left(\frac{1}{V^{2}}\right).$
We have $x_{t_{q}+1}=k$ under $y=k$, and Lemma 9 tells us that
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\right.$
×$\quad\left.\times\mathbf{1}\{t_{q}=s\}x_{t_{q}+1}\mathbf{1}\{z_{t}=k\}\mid y=k\right]$
$=\mathbb{E}\left[\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s\right]$
×$\quad\times\frac{P(t_{q}=s)}{T}w_{E}(k)$
=$=\begin{cases}\frac{P(t_{q}=s)}{T}\left(\frac{2}{V-1}+\frac{T-4}{(V-1)^{2}}\right)w_{E}(k)&\text{if $t\neq T$},\\ 0&\text{otherwise}.\end{cases}$
Next, we consider $x_{t}$. Note that $t=T$ forces $x_{t_{q}}$ to be $w_{E}(k)$ for the expectation to take non-zero value, which contradicts the input sequence condition. Hence we assume $t\neq T$ so that $z_{t_{q}}=z_{T}\neq k$.
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}x_{t}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s\right]$
$=\frac{1}{T}\left\{\mathbb{E}\left[\mathbf{1}\{z_{t_{q}}=k\}x_{t}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]\right.$
$\quad+\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}x_{t}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$\quad+\mathbb{E}\left[\mathbf{1}\{z_{t_{q}+1}=k\}x_{t}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$\quad+\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}x_{t}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$\quad+\left.\sum_{t^{\prime}}\mathbb{E}\left[\mathbf{1}\{z_{i}=k\}x_{t}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]\right\}$
$=\frac{1}{T}\left(\frac{2w_{E}(k)}{V-1}+\sum_{t^{\prime}}\frac{w_{E}(k)}{(V-1)^{2}}\right)$
$=\frac{1}{T}\cdot O\left(\frac{1}{V}\right).$
Lastly, regarding $x_{i}$ for $i\neq t,t_{q},t_{q}+1$, we also consider $t\neq T$ and it follows that
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid y=k,t_{q}=s\right]$
$=\frac{1}{T}\left\{\mathbb{E}\left[\mathbf{1}\{z_{t_{q}}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]\right.$
$\quad+\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$\quad+\mathbb{E}\left[\mathbf{1}\{z_{t_{q}+1}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$\quad+\mathbb{E}\left[\mathbf{1}\{z_{t}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]$
$\quad+\left.\sum_{t^{\prime}}\mathbb{E}\left[\mathbf{1}\{z_{i}=k\}x_{i}\mathbf{1}\{z_{t}=k\}\mid\cdot\right]\right\}$
$=\frac{1}{T}\left\{0+0+2\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{1}{V-1}\sum_{j\in\mathcal{V}\setminus\{q\}}\frac{w_{E}(j)}{(V-1)^{2}}\right.$
$\quad\left.+\sum_{t^{\prime}}\sum_{q\in\mathcal{V}\setminus\{k\}}\frac{1}{V-1}\sum_{j\in\mathcal{V}\setminus\{q\}}\frac{w_{E}(j)}{(V-1)^{3}}\right\}$
$=\frac{1}{T}\cdot O\left(\frac{1}{V}\right).$
Thus, we obtain Eq. 49. ∎
Lemma 14.
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. Then, we have
$\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\mid y=k\right]$
=$=\begin{cases}P(t_{q}=T-1)\left(r_{0}-\bar{r}_{1-T:0}\right)\\ \quad+P(t_{q}=T-1)\frac{T-2}{T}w_{E}(k)\\ \qquad+O\left(\frac{1}{V}\right)&\text{if $v=k$},\\ 0&\text{otherwise},\end{cases}$
where $q_{T,T}=(w_{E}(z_{T})+r_{0})-(\bar{w}_{E}(z_{1:T})+\bar{r}_{1-T:0})$.
Proof.
The inequality $v\neq k$ produces $\mathbf{1}\{z_{T}=k\}\mathbf{1}\{z_{T}=v\}=0$, and thus, we focus on $v=k$. Since $z_{T}$ is the second trigger token, we have $z_{t_{q}}=z_{T}=k$. This is achievable only when $t_{q}=T-1$.
$\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\mid y=k\right]$
$=\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}q_{T,T}\mid y=k\right]$
$=P(z_{T}=k\mid y=k)\mathbb{E}\left[q_{T,T}\mid y=k,z_{T}=k\right]$ (50)
$=\sum_{i=1}^{T}P(z_{T}=k,t_{q}=i\mid y=k)$ (51)
×$\quad\times\mathbb{E}\left[q_{T,T}\mid y=k,z_{T}=k\right]$
$=P(t_{q}=T-1\mid y=k)$
×$\quad\times P(z_{T}=k\mid y=k,t_{q}=T-1)$
×$\quad\times\mathbb{E}\left[q_{T,T}\mid y=k,z_{T}=k\right]$
$=P(t_{q}=T-1)\left\{w_{E}(k)+r_{0}-\bar{r}_{1-T:0}\right.$
$\quad\left.-\frac{1}{T}\left(2w_{E}(k)+\sum_{\begin{subarray}{c}1\leq j\leq V\\ j\neq k\end{subarray}}\frac{(T-2)}{V-1}w_{E}(j)\right)\right\}$
$=P(t_{q}=T-1)\left(r_{0}+\frac{T-2}{T}w_{E}(k)-\bar{r}_{1-T:0}\right)$
$\quad+O\left(\frac{1}{V}\right).$
Lemma 15.
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. Then, we have
$\mathbb{E}\left[\frac{1}{T}\sum_{t^{\prime}=1}^{T}\mathbf{1}\{z_{t^{\prime}}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\mid y=k\right]$
=$=\begin{cases}\frac{2P(t_{q}=T-1)}{T}\left\{r_{0}-\bar{r}_{1-T:0}\right\}\\ \quad+\frac{2P(t_{q}=T-1)}{T}\frac{T-2}{T}w_{E}(k)\\ \quad+\frac{1}{T}\cdot O\left(\frac{1}{V}\right)&\text{if $v=k$},\\ \frac{1}{T}O\left(\frac{1}{V}\right)&\text{otherwise}.\end{cases}$
where $q_{T,T}=(w_{E}(z_{T})+r_{0})-(\bar{w}_{E}(z_{1:T})+\bar{r}_{1-T:0})$.
Proof.
We first consider the case $v=k$. If the second trigger token $z_{T}=v=k$ and the output token is also $k$, it is impossible to have $t_{q}<T-1$ because it produces more than two trigger tokens in the input sequence. This means that we have, for $t^{\prime}\in[T-2]$, that
$\mathbb{E}\left[\mathbf{1}\{z_{t}^{\prime}=k\}\mathbf{1}\{z_{T}=k\}\mid y=k\right]=0.$
We can use Lemma 14 for $t^{\prime}=T$:
$\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\mid y=k\right]$
$=P(t_{q}=T-1)\left(r_{0}+\frac{T-2}{T}w_{E}(k)-\bar{r}_{1-T:0}\right)$
$\quad+O\left(\frac{1}{V}\right).$
Finally, for $t^{\prime}=T-1$, we note that the two conditions $y=k$ and $z_{T}=k$ imply $z_{T-1}=k$.
$\mathbb{E}\left[\mathbf{1}\{z_{T-1}=k\}q_{T,T}\mathbf{1}\{z_{T}=k\}\mid y=k\right]$
$=P(z_{T}=k\mid y=k)$
×$\quad\times\mathbb{E}\left[\mathbf{1}\{z_{T-1}=k\}q_{T,T}\mid y=k,z_{T}=k\right]$
$=P(z_{T}=k\mid y=k)\mathbb{E}\left[q_{T,T}\mid y=k,z_{T}=k\right]$
$=\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\mid y=k\right]$
$=P(t_{q}=T-1)\left(r_{0}+\frac{T-2}{T}w_{E}(k)-\bar{r}_{1-T:0}\right)$
$\quad+O\left(\frac{1}{V}\right),$
where the second-to-last expression follows from Eq. 50.
Now, we look at the case $v\neq k$. Different from $v=k$, we have
$\mathbb{E}\left[\mathbf{1}\{z_{T-1}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\mid y=k\right]=0,$
$\mathbb{E}\left[\mathbf{1}\{z_{T}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\mid y=k\right]=0.$
Furthermore, the trigger token is uniformly distributed, we have
$\mathbb{E}[\mathbf{1}\{z_{T}=v\}\mid y=k]$
$=P(z_{T}=v\mid y=k)$
$=\sum_{i=1}^{T}P(z_{T}=v,t_{q}=i\mid y=k)$
$=\sum_{i=1}^{T-2}P(z_{T}=v\mid y=k,t_{q}=i)$
×$\quad\times P(t_{q}=i\mid y=k)$
$=\frac{P(t_{q}\leq T-2)}{V-1}$
$=O\left(\frac{1}{V}\right).$
This leads to the following expression for $t^{\prime}\in[T-2]$
$\mathbb{E}\left[\mathbf{1}\{z_{t^{\prime}}=k\}q_{T,T}\mathbf{1}\{z_{T}=v\}\mid y=k\right]$
$=\mathbb{E}\left[\mathbf{1}\{z_{T}=v\}\right.$
×$\quad\left.\times\mathbb{E}\left[\mathbf{1}\{z_{t^{\prime}}=k\}q_{T,T}\mid y=k,z_{T}\right]\mid y=k\right]$
$=P(z_{T}=v\mid y=k)$
×$\quad\times\mathbb{E}\left[\mathbf{1}\{z_{t^{\prime}}=k\}q_{T,T}\mid y=k,z_{T}=v\right]$
$=O\left(\frac{1}{V}\right).$
This concludes the proof. ∎
C.2 global v.s. in-context
Proposition 4 (Restated).
Suppose a two-layer transformer is an associative memory transformer. Given a length-$T$ sequence $z_{1:T}$ where the last token is $z_{T}=q$, let the $t_{1}$-th token be $z_{t_{1}}=v_{1}$ following $z_{t_{1}-1}=q$, and let the $t_{2}$-th token be $z_{t_{2}}=v_{2}$ following $z_{t_{2}-1}=q$, where $v_{1}\neq v_{2}$ and $(3\leq)t_{1}<t_{2}$ without loss of generality. Also assume that there exist only one $v_{1}$ and $v_{2}$ in the context $z_{1:T}$. Then, the logit $\xi_{v_{1}}$ and $\xi_{v_{2}}$ can be expressed as follows:
$\xi_{v_{1}}$
$=\sigma\left(\frac{e}{t_{1}+e-1}\right)+\mathbf{1}\{v_{1}\in U_{q}\}\log{\pi_{b}(v_{1}\mid q)},$
$\xi_{v_{2}}$
$=\sigma\left(\frac{1+e}{t_{2}+e-1}\right)+\mathbf{1}\{v_{2}\in U_{q}\}\log{\pi_{b}(v_{2}\mid q)},$
where $\sigma(\cdot)$ represents the softmax function.
Proof.
We consider the output of the first layer of attention for $z_{t}$. Since we use the associative memory transformer, we substitute $W_{Q}^{1}=I$ and $W_{K}^{1}=\sum_{k\in Q}w_{E}(k)r_{-1}^{\top}$ into Eq. 1, we have
$x^{(1)}_{t}$
$=\Phi_{1}w_{E}(z_{1:t})\sigma((w_{E}(z_{1:t})+R_{1-t:0})^{\top}$
×$\qquad\times\left(\sum_{k\in Q}r_{-1}w_{E}(k)^{\top}\right)Iw_{E}(z_{t}))+w_{E}(z_{t})$
$=\Phi_{1}w_{E}(z_{1:t})\sigma\left(\begin{pmatrix}\left(w_{E}(z_{1})+r_{1-t}\right)^{\top}\\ \left(w_{E}(z_{2})+r_{2-t}\right)^{\top}\\ \vdots\\ \left(w_{E}(z_{t-1})+r_{-1}\right)^{\top}\\ \left(w_{E}(z_{t})+r_{0}\right)^{\top}\end{pmatrix}\right.$
×$\qquad\qquad\left.\times r_{-1}w_{E}(z_{t})^{\top}w_{E}(z_{t})\right)+w_{E}(z_{t})$
$=\Phi_{1}w_{E}(z_{1:t})\sigma\left(\begin{pmatrix}0\\ 0\\ \vdots\\ 1\\ 0\end{pmatrix}\right)+w_{E}(z_{t})$ (52)
$=w_{E}(z_{t})+\frac{e(\Phi_{1}w_{E}(z_{t-1}))}{t+e-1}+\sum_{\begin{subarray}{c}1\leq i\leq t\\ i\neq t-1\end{subarray}}\frac{\Phi_{1}w_{E}(z_{i})}{t+e-1},$ (53)
where $e$ is Euler’s number. In the transformation above, Eq. 52 follows from the near-orthogonality among the relative positional encodings and the embedding vectors Next, we calculate the output of the second layer attention block. We first consider softmax values and then attention head $W_{O}^{2}W_{V}^{2}$. Similarly to the first attention layer, we use Eq. 53 and ideal matrices $W_{Q}^{2}=I$ and $W_{K}^{2}=\sum_{k\in Q}w_{E}(k)(\Phi_{1}w_{E}(k))^{\top}$ to calculate softmax. For any $t$, we can express $(x_{1:t}^{(1)})^{\top}(W_{K}^{2})^{\top}W_{Q}^{2}x_{t}^{(1)}$ as
$\begin{pmatrix}\left(w_{E}(z_{1})+\Phi_{1}w_{E}(z_{1})\right)^{\top}\\ \left(w_{E}(z_{2})+\frac{e(\Phi_{1}w_{E}(z_{1}))}{e+1}+\sum\frac{\Phi_{1}w_{E}(z_{i})}{e+1}\right)^{\top}\\ \vdots\\ \left(w_{E}(z_{t-1})+\frac{e(\Phi_{1}w_{E}(z_{t-2}))}{t+e-2}+\sum\frac{\Phi_{1}w_{E}(z_{i})}{t+e-2}\right)^{\top}\\ \left(w_{E}(z_{t})+\frac{e(\Phi_{1}w_{E}(z_{t-1}))}{t+e-1}+\sum\frac{\Phi_{1}w_{E}(z_{i})}{t+e-1}\right)^{\top}\end{pmatrix}$
×$\left.\times\left(\sum_{k\in Q}(\Phi_{1}w_{E}(k))w_{E}(k)^{\top}\right)Ix_{t}^{(1)}\right),$
where the summation in the $t$-th row is given by $\sum_{\begin{subarray}{c}1\leq i\leq t\\ i\neq t-1\end{subarray}}$. Since $x_{T}^{(1)}=w_{E}(q)+\frac{e(\Phi_{1}w_{E}(z_{T-1}))}{T+e-1}+\sum_{\begin{subarray}{c}1\leq i\leq T\\ i\neq T-1\end{subarray}}\frac{\Phi_{1}w_{E}(z_{i})}{T+e-1}$, and thanks to the near-orthogonality, we have
$\sigma\left((x_{1:T}^{(1)})^{\top}(W_{K}^{2})^{\top}W_{Q}^{2}x_{T}^{(1)}\right)$
$=\sigma\left(p\right),$
where each entry of $p\in\mathbb{R}^{T}$ is
$p_{t}=\begin{cases}0&\text{if $t<t_{1}-1$},\\ 1/(t_{1}+e-2)&\text{if $t=t_{1}-1$},\\ e/(t_{1}+e-1)&\text{if $t=t_{1}$},\\ 1/(i+e-1)&\text{if $t_{1}<t<t_{2}-1$},\\ 2/(t_{2}+e-2)&\text{if $t=t_{2}-1$},\\ (1+e)/(t_{2}+e-1)&\text{if $t=t_{2}$},\\ 2/(i+e-1)&\text{if $t_{2}<t<T$},\\ 3/(T+e-1)&\text{if $t=T$}.\end{cases}$
Back to Eq. 2, we can now write $x_{T}^{(2)}$ with $W_{O}^{2}=\sum_{v=1}^{V}w_{U}(v)(W_{V}^{2}w_{E}(v))^{\top}$
$x_{T}^{(2)}$
$=\sum_{t=1}^{T}W_{O}^{2}W_{V}^{2}\sigma(p_{t})x_{t}^{(1)}+x_{T}^{(1)}$
$=\sum_{t=1}^{T}\sigma(p_{t})\sum_{v=1}^{V}w_{U}(v)(W_{V}^{2}w_{E}(v))^{\top}$
×$\quad\times W_{V}^{2}\left(w_{E}(z_{t})+\frac{e(\Phi_{1}w_{E}(z_{t-1}))}{t+e-1}\right.$
$\qquad\left.+\sum_{\begin{subarray}{c}1\leq i\leq t\\ i\neq t-1\end{subarray}}\frac{\Phi_{1}w_{E}(z_{i})}{t+e-1}\right)+x_{T}^{(2)}$
$=\sum_{t=1}^{T}\sigma(p_{t})w_{U}(z_{t})+w_{E}(q)$
$\quad+\frac{e(\Phi_{1}w_{E}(z_{T-1}))}{T+e-1}+\sum_{\begin{subarray}{c}1\leq i\leq t\\ i\neq T-1\end{subarray}}\frac{\Phi_{1}w_{E}(z_{i})}{T+e-1}$
Finally, the feed-forward network in the second layer is applied to $x_{T}^{(2)}$, and we obtain the following:
$x_{T}$ $=W_{F}x_{T}^{(2)}+x_{T}^{(2)}$
$=\left(\sum_{v=1}^{V}\sum_{u\in U_{v}}\log{\pi_{b}(u\mid v)}w_{U}(u)w_{E}(v)^{\top}\right)$
×$\quad\times\left(w_{E}(q)+\sum_{t=1}^{T}\sigma(p_{t})w_{U}(z_{t})+C\right)+x_{T}^{(2)}$
$=\sum_{u\in U_{q}}\log{\pi_{b}(u\mid q)}w_{U}(u)+x_{T}^{(2)}$
$=\sum_{u\in U_{q}}\log{\pi_{b}(u\mid q)}w_{U}(u)$
$\quad+\sum_{t=1}^{T}\sigma(p_{t})w_{U}(z_{t})+w_{E}(q)+C,$
where $C=\frac{e(\Phi_{1}w_{E}(z_{T-1}))}{T+e-1}+\sum_{\begin{subarray}{c}1\leq i\leq t\\ i\neq T-1\end{subarray}}\frac{\Phi_{1}w_{E}(z_{i})}{T+e-1}$.
We can compute the logits $\xi_{v}$ for $v\in\mathcal{V}$ after the unembedding layer. Since the row vectors of $W_{U}$ are also near-orthogonal to each other, it is sufficient to look at the coefficients of $w_{U}(v_{1})$ and $w_{U}(v_{2})$. Since $v_{1}$ and $v_{2}$ appear only once at $t=t_{1}$ and $t=t_{2}$ respectively in the context $z_{1:T}$, the logits are calculated as follows:
$\xi_{v_{1}}$ $=\sigma\left(\frac{e}{t_{1}+e-1}\right)+\mathbf{1}\{v_{1}\in U_{q}\}\log{\pi_{b}(v_{1}\mid q)},$
$\xi_{v_{2}}$ $=\sigma\left(\frac{1+e}{t_{2}+e-1}\right)+\mathbf{1}\{v_{2}\in U_{q}\}\log{\pi_{b}(v_{2}\mid q)}.$
Proposition 5 (Restated).
Suppose a two-layer transformer is a stronger associative memory transformer as in Eqs. 10, 11, and 12. Given a length-$T$ sequence $z_{1:T}$ where the last token is $z_{T}=q$, let $f(v)$ be the number of token pattern "$qv$" appearing in the sequence for vocabulary $v\in\mathcal{V}$. For large enough $\tau_{1}$ and $\tau_{2}$, the logits $\xi_{v}$ can be expressed as follows:
$\xi_{v}$ $\underset{\begin{subarray}{c}\tau_{1}\\ \tau_{2}\end{subarray}}{\approx}\log{\pi_{b}(v\mid q)}$
$\quad+\tau_{3}\cdot\frac{f(v)+\mathbf{1}\{v=q\}\mathbf{1}\{z_{1}=q\}}{\left(\sum_{v^{\prime}=1}^{V}f(v^{\prime})\right)+\mathbf{1}\{z_{1}=q\}}.$ (54)
Proof.
The proof proceeds in the same manner as in Prop. 2. With the general associative memory $W_{K}^{1}$, the attention values for $z_{t}$ is expressed as
$x^{(1)}_{t}$
$=\Phi_{1}w_{E}(z_{1:t})\sigma((w_{E}(z_{1:t})+R_{1-t:0})^{\top}$
×$\quad\times\left(\tau_{1}\sum_{k\in Q}r_{-1}w_{E}(k)^{\top}\right)Iw_{E}(z_{t}))+w_{E}(z_{t})$
$=\Phi_{1}w_{E}(z_{1:t})\sigma\left(\tau_{1}\begin{pmatrix}\left(w_{E}(z_{1})+r_{1-t}\right)^{\top}\\ \left(w_{E}(z_{2})+r_{2-t}\right)^{\top}\\ \vdots\\ \left(w_{E}(z_{t-1})+r_{-1}\right)^{\top}\\ \left(w_{E}(z_{t})+r_{0}\right)^{\top}\end{pmatrix}\right.$
×$\qquad\qquad\left.\times r_{-1}w_{E}(z_{t})^{\top}w_{E}(z_{t})\right)+w_{E}(z_{t})$
$=\Phi_{1}w_{E}(z_{1:t})\sigma\left(\begin{pmatrix}0\\ 0\\ \vdots\\ \tau_{1}\\ 0\end{pmatrix}\right)+w_{E}(z_{t})$
$\underset{\tau_{1}}{\approx}\Phi_{1}w_{E}(z_{t-1})+w_{E}(z_{t}).$
Now the second layer attention has the form
$\sigma\left((x_{1:t}^{(1)})^{\top}(W_{K}^{2})^{\top}W_{Q}^{2}x_{t}^{(1)}\right)$
$\underset{\tau_{1}}{\approx}$ $\sigma\left(\begin{pmatrix}\left(w_{E}(z_{1})+\Phi_{1}w_{E}(z_{1})\right)^{\top}\\ \left(w_{E}(z_{2})+\Phi_{1}w_{E}(z_{1})\right)^{\top}\\ \vdots\\ \left(w_{E}(z_{t-1})+\Phi_{1}w_{E}(z_{t-2})\right)^{\top}\\ \left(w_{E}(z_{t})+\Phi_{1}w_{E}(z_{t-1})\right)^{\top}\end{pmatrix}\right.$
×$\qquad\left.\times\left(\tau_{2}\sum_{k\in Q}(\Phi_{1}w_{E}(k))w_{E}(k)^{\top}\right)Ix_{t}^{(1)}\right).$
We note that $x_{T}^{(1)}=w_{E}(q)+\Phi_{1}w_{E}(z_{T-1})$, and consider the case $t=T$ in the above equation. Then, we will have
$\sigma\left((x_{1:T}^{(1)})^{\top}(W_{K}^{2})^{\top}W_{Q}^{2}x_{T}^{(1)}\right)$
$\underset{\tau_{1}}{\approx}$ $\sigma\left(\begin{pmatrix}\left(w_{E}(z_{1})+\Phi_{1}w_{E}(z_{1})\right)^{\top}\\ \left(w_{E}(z_{2})+\Phi_{1}w_{E}(z_{1})\right)^{\top}\\ \vdots\\ \left(w_{E}(z_{T-1})+\Phi_{1}w_{E}(z_{T-2})\right)^{\top}\\ \left(w_{E}(z_{T})+\Phi_{1}w_{E}(z_{T-1})\right)^{\top}\end{pmatrix}\right.$
×$\qquad\left.\times\left(\tau_{2}\sum_{k\in Q}(\Phi_{1}w_{E}(k))w_{E}(k)^{\top}\right)Ix_{T}^{(1)}\right)$
$=$ $\sigma\left(\tau_{2}\begin{pmatrix}\left(w_{E}(z_{1})+\Phi_{1}w_{E}(z_{1})\right)^{\top}\\ \left(w_{E}(z_{2})+\Phi_{1}w_{E}(z_{1})\right)^{\top}\\ \vdots\\ \left(w_{E}(a)+\Phi_{1}w_{E}(q)\right)^{\top}\\ \vdots\\ \left(w_{E}(b)+\Phi_{1}w_{E}(q)\right)^{\top}\\ \vdots\\ \left(w_{E}(z_{T})+\Phi_{1}w_{E}(z_{T-1})\right)^{\top}\end{pmatrix}\Phi_{1}w_{E}(q)\right)$
$=$ $\sigma\left(\sum_{i\in I_{q}\cup I_{1}}\tau_{2}\mathbf{e}_{i}\right),$
where $I_{q}=\{t\mid x_{t}^{(1)}=w_{E}(v^{\prime})+\Phi_{1}w_{E}(q),\exists v^{\prime}\in\mathcal{V}\}$ is the index set of $x_{t}^{(1)}$ containing $\Phi_{1}w_{E}(q)$, and $I_{1}=\{1\}$ if $z_{1}=q$ and $I_{1}=\emptyset$ otherwise. Also, $\mathbf{e}_{i}$ denote the unit vector where only the $i$-th component is 1, and all other components are 0. Note that $|I_{q}\cup I_{1}|=\left(\sum_{v=1}^{V}f(v)\right)+\mathbf{1}\{z_{1}=q\}$.
Going back to the output of the whole attention block including $W_{O}^{2}$ and $W_{V}^{2}$, we have from Eq. 2 that
$x_{T}^{(2)}$ $\underset{\tau_{1}}{\approx}\sum_{t=1}^{T}W_{O}^{2}W_{V}^{2}\sigma\big{(}\sum_{i\in I_{q}\cup I_{1}}\tau_{2}\mathbf{e}_{i}\big{)}_{t}x_{t}^{(1)}+x_{T}^{(1)}$
$\underset{\begin{subarray}{c}\tau_{1}\\ \tau_{2}\end{subarray}}{\approx}\frac{1}{|I_{q}\cup I_{1}|}\sum_{i\in I_{q}\cup I_{1}}W_{O}^{2}W_{V}^{2}x_{i}^{(1)}+x_{T}^{(1)}$
$=\frac{1}{|I_{q}\cup I_{1}|}\sum_{v=1}^{V}\tau_{3}w_{U}(v)(W_{V}^{2}w_{E}(v))^{\top}$
×$\quad\times\left\{f(v)\left(W_{V}^{2}(w_{E}(v)+\Phi_{1}w_{E}(q))\right)\right\}$
$\quad+\mathbf{1}\{z_{1}=q\}\left(W_{V}^{2}(w_{E}(q)+\Phi_{1}w_{E}(q))\right)$
$\quad+x_{T}^{(1)}$
$=\sum_{v=1}^{V}\frac{f(v)+\mathbf{1}\{v=q\}\mathbf{1}\{z_{1}=q\}}{\left(\sum_{v^{\prime}=1}^{V}f(v^{\prime})\right)+\mathbf{1}\{z_{1}=q\}}$
×$\quad\times\tau_{3}w_{U}(v)+w_{E}(q)+\Phi_{1}w_{E}(z_{T-1})$
$:=\sum_{v=1}^{V}g(v)\cdot w_{U}(v)$
$\quad+w_{E}(q)+\Phi_{1}w_{E}(z_{T-1}).$
Lastly, through the feed-forward network $W_{F}$, we obtain the outcome.
$x_{T}$
$=W_{F}x_{T}^{(2)}+x_{T}^{(2)}$
$\underset{\begin{subarray}{c}\tau_{1}\\ \tau_{2}\end{subarray}}{\approx}\left(\sum_{v=1}^{V}\sum_{u\in U_{v}}\log{\pi_{b}(u\mid v)}w_{U}(u)w_{E}(v)^{\top}\right)$
×$\quad\times\left(w_{E}(q)+\Phi_{1}w_{E}(z_{T-1})+\sum_{v=1}^{V}g(v)w_{U}(v)\right)$
$\quad+x_{T}^{(2)}$
$=\sum_{u\in U_{q}}\log{\pi_{b}(u\mid q)}w_{U}(u)$
$\quad+\left(\sum_{v=1}^{V}g(v)w_{U}(v)\right)+w_{E}(q)+\Phi_{1}w_{E}(z_{T-1})$
$=\sum_{v=1}^{V}\left(\mathbf{1}\{v\in U_{q}\}\log{\pi_{b}(v\mid q)}\right.$
$\qquad\left.+\tau_{3}\cdot\frac{f(v)+\mathbf{1}\{v=q\}\mathbf{1}\{z_{1}=q\}}{\left(\sum_{v^{\prime}=1}^{V}f(v^{\prime})\right)+\mathbf{1}\{z_{1}=q\}}\right)\cdot w_{U}(v)$
$\quad+w_{E}(q)+\Phi_{1}w_{E}(z_{T-1}).$
The proof concludes by examining the coefficients of $w_{U}(v)$. ∎
C.3 Induction head without positional encoding
In this section, we provide the construction of a three-layer transformer with no positional encoding that achieves the induction head mechanism.
We first introduce the following theorem[27] stating that transformer can implement APE using one transformer block.
Theorem 4 (Theorem 1, [27]).
Let $z_{1:T+1}$ be an input sequence of length $T+1$, where $z_{1}=\langle bos\rangle$. Then, there exist a transformer block consisting of an attention block and feed-froward network that recovers absolute positions $[1,2,\dots T+1]$ and writes in the hidden state $H^{(1)}$.
In the proof, they prepare extra $3$ dimensions for word embeddings to effectively reconstruct the positional information:
$W_{E}=\begin{pmatrix}1&1&1&\cdots&1\\ 1&0&0&\cdots&0\\ 0&0&0&\cdots&0\\ a^{e}_{1,1}&a^{e}_{1,2}&a^{e}_{1,3}&\cdots&a^{e}_{1,V}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a^{e}_{d,1}&a^{e}_{d,2}&a^{e}_{d,3}&\cdots&a^{e}_{d,V}\end{pmatrix},$
where $\langle bos\rangle$ is represented by the first column without loss of generality.
Theorem 4 states that the hidden state $H^{(1)}$ has the following form after the appropriate transformer block:
$H^{(1)}=\begin{pmatrix}1&1&1&\cdots&1\\ 1&0&0&\cdots&0\\ 1&2&3&\cdots&T+1\\ a_{1,1}&a_{1,2}&a_{1,3}&\cdots&a_{1,T+1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{d,1}&a_{d,2}&a_{d,3}&\cdots&a_{d,T+1}\end{pmatrix},$
Now we show that a three-layer transformer with no positional encoding can implement induction head mechanism.
Proposition 6 (Restated).
There exists construction of a three-layer transformer without positional encoding that achieves induction head mechanism.
Proof.
The first layer of the transformer block is given by the construction in Theorem 4, which gives us the hidden state:
$H^{(1)}=\begin{pmatrix}1&1&1&\cdots&1\\ 1&0&0&\cdots&0\\ 1&2&3&\cdots&T+1\\ a_{1,1}&a_{1,2}&a_{1,3}&\cdots&a_{1,T+1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{d,1}&a_{d,2}&a_{d,3}&\cdots&a_{d,T+1}\end{pmatrix},$
where each entry denoted by $a_{*,*}$ is a Gaussian random value to guarantee near-orthogonality.
Then, for the second transformer block, we use the following matrix weights.
$W_{Q}^{2}$ $=\begin{pmatrix}1&0&0&\cdots&0\\ 0&0&0&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&0\end{pmatrix},$
$W_{K}^{2}$ $=\begin{pmatrix}0&0&C&\cdots&0\\ 0&0&0&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&0\end{pmatrix},$
$W_{V}^{2}$ $=\begin{pmatrix}0&0&0&0&\cdots&0\\ 0&0&0&0&\cdots&0\\ 0&0&0&0&\cdots&0\\ 0&0&0&w^{v}_{1,1}&\cdots&w^{v}_{1,d}\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&w^{v}_{d,1}&\cdots&w^{v}_{d,d}\end{pmatrix},$
$W_{O}^{2}$ $=\begin{pmatrix}0&0&0&0&\cdots&0\\ 0&0&0&0&\cdots&0\\ 0&0&0&0&\cdots&0\\ 0&0&0&w^{o}_{1,1}&\cdots&w^{o}_{1,d}\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&w^{o}_{d,1}&\cdots&w^{o}_{d,d}\end{pmatrix},$
where $C\in\mathbb{R}_{+}$ is a constant value, and $w^{*}_{*,*}$ denotes the Gaussian entry to guarantee the near-orthogonality. With these matrix, the query and key vectors are represented by
$q_{t}$ $=(1,0,\dots,0)^{\top},$
$k_{i}$ $=(C\cdot i,0,\dots 0)^{\top}.$
Thus, the block outputs
$\sum_{i=1}^{t-1}\sigma(\langle k_{i},q_{t}\rangle)W_{O}^{2}W_{V}^{2}h^{(1)}_{i}$ (55)
$\approx W_{O}^{2}W_{V}^{2}h_{t-1}^{(1)},$ (56)
where the approximation holds for sufficiently large $C$ and $h_{t-1}^{(1)}$ is the $t-1$-th column of $H^{(1)}$. Here, we assume that the attention block uses strict causal attention, which attends only to positions in $\{1,2,\dots,t-1\}$ for the token at position $t$.
Eq. 56 implies that the attention block attends only to the previous token, thus forming a previous token head. With the residual connection, the hidden state $H^{(2)}$ is represented as:
$H^{(2)}$
$=\begin{pmatrix}1&1&1&\cdots&1\\ 1&0&0&\cdots&0\\ 1&2&3&\cdots&T+1\\ a_{1,1}&a_{1,2}&a_{1,3}&\cdots&a_{1,T+1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{d,1}&a_{d,2}&a_{d,3}&\cdots&a_{d,T+1}\end{pmatrix}$
$\quad+\begin{pmatrix}0&0&0&\cdots&0\\ 0&0&0&\cdots&0\\ 0&0&0&\cdots&0\\ \mathbf{0}&\Phi_{2}\mathbf{a}_{1}&\Phi_{2}\mathbf{a}_{2}&\cdots&\Phi_{2}\mathbf{a}_{T}\end{pmatrix},$
where $\mathbf{a}_{i}=(a_{1,i},\dots,a_{d,i})^{\top}$ and $\Phi_{2}$ is the bottom right sub-matrix from $(W_{O}^{2}W_{V}^{2})_{4:d+3,4:d+3}$.
Finally, in the third transformer block, we set the following attention matrix weights:
$W_{Q}^{3}$ $=\begin{pmatrix}0&0&0&0&0&\cdots&0\\ 0&0&0&0&0&\cdots&0\\ 0&0&0&0&0&\cdots&0\\ 0&0&0&1&0&\cdots&0\\ 0&0&0&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&0&\cdots&1\end{pmatrix},$
$W_{K}^{3}$ $=\begin{pmatrix}0&0&0&0&\cdots&0\\ 0&0&0&0&\cdots&0\\ 0&0&0&0&\cdots&0\\ 0&0&0&w^{k^{\prime}}_{1,1}&\cdots&w^{k^{\prime}}_{1,d}\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&w^{k^{\prime}}_{d,1}&\cdots&w^{k^{\prime}}_{d,d}\end{pmatrix},$
$W_{V}^{3}$ $=\begin{pmatrix}0&0&0&0&\cdots&0\\ 0&0&0&0&\cdots&0\\ 0&0&0&0&\cdots&0\\ 0&0&0&w^{v^{\prime}}_{1,1}&\cdots&w^{v^{\prime}}_{1,d}\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&w^{v^{\prime}}_{d,1}&\cdots&w^{v^{\prime}}_{d,d}\end{pmatrix},$
$W_{O}^{3}$ $=\begin{pmatrix}0&0&0&0&\cdots&0\\ 0&0&0&0&\cdots&0\\ 0&0&0&0&\cdots&0\\ 0&0&0&w^{o^{\prime}}_{1,1}&\cdots&w^{o^{\prime}}_{1,d}\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&w^{o^{\prime}}_{d,1}&\cdots&w^{o^{\prime}}_{d,d}\end{pmatrix},$
where $w^{v^{\prime}}_{*,*}$ denotes the Gaussian entry to guarantee the near-orthogonality, and the sub-matrices in the bottom right of $W_{K}^{3}$ and $W_{O}^{3}$ are constructed the same way as $W_{K}^{2}$ and $W_{O}^{2}$ in Lemma 1. The result of this calculation is the same as in Lemma 1 except for the first three rows. Thus, the induction head mechanism is implemented in the three-layer transformer with no positional encoding. ∎
Appendix D Experimental setup
D.1 Oversight of in-context knowledge
We conducted our experiments using the following setup, carefully designed to evaluate the performance of a transformer model with APE and RPE.
Sequence generation
Following Bietti et al. [4], we define $\pi_{u}$ and $\pi_{b}$ as unigram and bigram character-level distributions estimated from the Tiny Shakespeare dataset [26]. We sample triggers from $\pi_{q}=\pi_{u}$, and corresponding outputs $o_{k}$ are also sampled uniformly for each sequence generation. The vocabulary size $|\mathcal{V}|$ is $65$.
Data Arguments
The model was trained on sequences with length of 256 tokens for the previous token head experiment and 128 tokens for the generalization experiment. The number of trigger tokens was set to 5, and we did not fix trigger tokens, i.e., all trigger tokens were sampled uniformly.
Model Configuration
The transformer model used for the experiments had the following configuration:
  • Model dimension: 256
  • Vocabulary size: 65
  • Input sequences were capped at a maximum length of 256 tokens.
  • The embedding layer $W_{E}$ and unembedding layer $W_{U}$ were frozen during training.
  • The model did not use $W_{F}$ in the second layer to focus on the role of induction head.
  • The model has $0.5$M parameters.
Optimization Setup
The optimization process was conducted using the following hyperparameters:
  • Optimizer: Stochastic Gradient Descent (SGD) with momentum
  • Batch size: 512
  • Learning rate: 0.2
  • Momentum: 0.9
  • Weight decay: 0.0001
Training Strategy
  • The model was trained for a total of 1000 iterations, which takes 0.5 hours of A100 GPU time.
  • Loss computation was restricted to output tokens.
D.2 Global vs in-context
We conducted our experiments using the following setup for the evaluation of performance of a transformer model with RPE trained on The Google analogy dataset.
Data Arguments
The model was trained on sequences with a maximum length of 256 tokens.
Since The Google analogy dataset offers analogy pairs, such as (Tokyo, Japan), the bigram conditionals will mostly be in the form like $\pi_{b}(\text{Japan}\mid\text{Tokyo})=1$. We constructed new bigram conditionals by adding some randomenss.
We first sample 10 fake target words, such as USA, China for each source word. Then, the conditionals will be calculated as follows:
  1. 1.
    count the number $c_{A}(B)$ of analogical pair $(A,B)$.
  2. 2.
    sample $p_{A}\in[0.01,0.1]$ for each source word $A$.
  3. 3.
    consider the analogical pair $(A,B_{i})$ appear $p_{r}c_{A}(B)$ times, where $B_{i}$ is the fake target words for source word $A$.
  4. 4.
    calculate the bigram conditionals based on the number of appearance of analogical pairs.
Everytime we generate an input sequence, we sample 5 trigger tokens from the set of all source words, and uniformly selects the corresponding output tokens from all vocabulary. we start from a source word that is sampled uniformly from the set of source words. Then, we transition with either the bigram conditionals or to the output token if the current token is the trigger token. Then we place comma, and sample uniformly a new word from the set of source words for the next token, and repeat this process until the maximum length is achieved.
Model Configuration
The transformer model used for the experiments had the following architecture and configuration:
  • Model dimension: 512
  • Vocabulary size: 233
  • Input sequences were capped at a maximum length of 256 tokens.
  • The embedding and output layers were frozen during training.
  • The model has $2$M parameters.
Optimization Setup
The optimization process was conducted using the following hyperparameters:
  • Optimizer: Stochastic Gradient Descent (SGD) with momentum
  • Batch size: 512
  • Learning rate: 0.1
  • Momentum: 0.9
  • Weight decay: 0.0001
Training Strategy
  • The model was trained for a total of 100,000 iterations, which required 9 hours of A100 GPU time.
  • Loss computation was applied to all tokens except for commas.
D.3 About datasets
The Tiny Shakespeare dataset is licensed under the Apache License, Version 2.0, and we can freely use the content as long as we comply with the terms of the license. Similarly, the Google Analogy dataset is licensed under CC BY-NC 4.0, allowing use of the content for non-commercial purposes with appropriate attribution. Our use of these datasets is consistent with their intended purposes. Furthermore, there is no explicit measure that should be taken to check for personal information or offensive content.
Appendix E Other experimental results and analyses
(a) Attention patterns of $\operatorname{TF}_{\text{APE}}$
(b) Attention patterns of $\operatorname{TF}_{\text{RPE}}$
Figure 4: Two-layer transformers with APE and RPE that are trained on sequences of length $128$ show different first-layer attention patterns for length-$256$ sequences. The previous token head of $\operatorname{TF}_{\text{APE}}$ do not function well for positions $t>100$, while $\operatorname{TF}_{\text{RPE}}$ attends to previous tokens regardless of the positions.
E.1 Generalization to longer sequence
acc. ($t<128$) acc. ($t<256$) score ($t<128$) score ($t<256$)
$\operatorname{TF}_{\text{APE}}$ $0.8278\pm 0.0000$ $0.7925\pm 0.0000$ $0.5460\pm 0.0007$ $0.1979\pm 0.0001$
$\operatorname{TF}_{\text{RPE}}$ $0.8727\pm 0.0000$ $0.8974\pm 0.0000$ $0.9513\pm 0.0000$ $0.9337\pm 0.0001$
Table 3: The evaluation results under the change in sequence length from $128$ during training to $256$ during evaluation. Transformer with APE degrades its accuracy and attention score as the sequence length grows, while transformer with RPE shows a stable accuracy and score.
To evaluate the generalization ability of $\operatorname{TF}_{\text{ape}}$ and $\operatorname{TF}_{\text{rpe}}$, we trained these two types of transformers with sequences of length $128$ generated from the bigram model. Details regarding the training procedure are provided in the Appendix D. We measured the accuracy of $\operatorname{TF}_{\text{APE}}$ and $\operatorname{TF}_{\text{RPE}}$ with respect to the second and subsequent occurrences of the output tokens. In addition, We present the memory recall to examine the generalization.
The results are shown in Tab. 3. In terms of accuracy, transformer with RPE raises its accuracy even when the input sequence length increased to $256$, while transformer with APE dropped its accuracy to $79.25\%$. Although the difference may seem trivial, it is worth noting that the accuracy for $t<256$ is computed through the whole sequence including $t<128$. This means that transformer with APE is more likely to fail to predict output tokens at positions $256128<t<$ 256. Looking at the attention scores, we can confirm that $\operatorname{TF}_{\text{APE}}$ is incapable of attending to previous tokens, even for positions $t<128$. In contrast, $\operatorname{TF}_{\text{RPE}}$ is trained to have a previous attention head that keeps its performance no matter how long the sequence is.
In fact, the attention pattern heatmaps in Fig. 4 in Appendix E tell us that the attention is no longer directed to the previous token after $t>100$ for $\operatorname{TF}_{\text{APE}}$, while $\operatorname{TF}_{\text{RPE}}$ continues to attend to the previous token. These experimental results match our theoretical result, where RPE is better suited for length generalization. It can be seen from Fig. 4(b) that the first-layer attention head of $\operatorname{TF}_{\text{RPE}}$ sometimes attends to the current token. This phenomenon is also explainable from the proof of Theorem 3, which shows that the attention matrix $W_{K}^{1}$ not only learns the associations between $w_{E}(v)$ and $r_{-1}$, but also the pair of $w_{E}(v)$ and $r_{0}$.
E.2 Discussion of collision of context information
We discuss the result shown in Fig. 3. The figure illustrates that it is more likely to output global knowledge when the number of $B_{1}$ and $B_{2}$ are the same, or when the number of $B_{1}$ or $B_{2}$ is almost maximum. Using Proposition 2, we can explain these phenomena by considering inequalities using the global knowledge $\log{\pi_{b}(\cdot\mid A)}$ and the strength of in-context knowledge $\tau_{3}$. Using Proposition 2, we can explain these phenomena as follows. If we have the relationships $\log{\pi_{b}(B_{*}\mid A)}+\tau_{3}>\log{\pi_{b}(B\mid A)}>\log{\pi_{b}(B_{*}\mid A)}+\tau_{3}/2$, where $B_{*}$ represents $B_{1}$ and $B_{2}$, the model outputs global knowledge $B$ when $AB_{1}$ and $AB_{2}$ occur equally frequently in the prompt. On the other hand, if we have $\log{\pi_{b}(B_{2}\mid A)}+\tau_{3}>\log{\pi_{b}(B\mid A)}>\log{\pi_{b}(B_{1}\mid A)}+\tau_{3}$ and $\log{\pi_{b}(B\mid A)}>\log{\pi_{b}(B_{2}\mid A)}$, the prediction of the model changes to $B_{2}$ to $B$ as the number of $B_{1}$ increases.